### v1

# AutoMIR: Effective Zero-Shot Medical Information Retrieval without Relevance Labels

Abstract.
Medical information retrieval (MIR) is essential for retrieving relevant medical knowledge from diverse sources, including electronic health records, scientific literature, and medical databases. However, achieving effective zero-shot dense retrieval in the medical domain poses substantial challenges due to the lack of relevance-labeled data. In this paper, we introduce a novel approach called Self-Learning Hypothetical Document Embeddings (SL-HyDE) to tackle this issue. SL-HyDE leverages large language models (LLMs) as generators to generate hypothetical documents based on a given query. These generated documents encapsulate key medical context, guiding a dense retriever in identifying the most relevant documents. The self-learning framework progressively refines both pseudo-document generation and retrieval, utilizing unlabeled medical corpora without requiring any relevance-labeled data. Additionally, we present the Chinese Medical Information Retrieval Benchmark (CMIRB), a comprehensive evaluation framework grounded in real-world medical scenarios, encompassing five tasks and ten datasets. By benchmarking ten models on CMIRB, we establish a rigorous standard for evaluating medical information retrieval systems. Experimental results demonstrate that SL-HyDE significantly surpasses existing methods in retrieval accuracy while showcasing strong generalization and scalability across various LLM and retriever configurations. CMIRB data and evaluation code are publicly available at: https://github.com/CMIRB-benchmark/CMIRB
Medical Information Retrieval, Large Language Models, Zero-shot Retrieval
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Retrieval models and ranking
1. INTRODUCTION
Medical information retrieval (MIR) (Luo et al., 2008; Goeuriot et al., 2016) focuses on retrieving relevant medical information from extensive data sources, such as electronic health records, scientific papers, and medical knowledge databases, based on specific medical queries. Its applications are wide-ranging, supporting doctors in clinical decision-making (Sivarajkumar et al., 2024), assisting patients in finding health information (McGowan et al., 2009), and aiding researchers in accessing relevant studies (Zheng and Yu, 2015).
In information retrieval (IR), there are two mainstream paradigms: lexical-based sparse retrieval methods, such as BM25 (Robertson et al., 2009), and embedding-based dense retrieval approaches (Xiong et al., 2020; Ding et al., 2020; Izacard et al., 2021). Dense retrievers have shown strong performance, particularly in domains where large annotated datasets are available (Karpukhin et al., 2020). These datasets are typically created by manually matching queries with relevant documents, a process that becomes impractical when dealing with collections containing millions or billions of items. As a result, many retrieval tasks face the challenge of having limited or no in-domain labeled examples.
Given these challenges, several studies (Karpukhin et al., 2020; Xiong et al., 2020) have explored transfer learning for zero-shot retrieval, where dense retrievers are trained on large, high-resource datasets and evaluated on queries from different domains. MS MARCO (Bajaj et al., 2016), a dataset containing a substantial number of manually annotated query-document pairs, is commonly used for this purpose. An alternative to transfer learning is contrastive learning (Gao et al., 2021), which follows a two-stage pretraining and fine-tuning pipeline to develop general-purpose text embedding models. The pretraining stage leverages weakly supervised data gathered through large-scale web crawling, while the fine-tuning stage uses high-quality text pairs derived from data mining or manual annotation (Xiao et al., 2024; Li et al., 2023). However, the availability of such large-scale datasets or high-quality text pairs cannot always be assumed, particularly in non-English languages or specialized domains.
Recently, large language models (LLMs) have demonstrated exceptional performance in zero-resource retrieval scenarios, primarily due to their extensive knowledge and robust text generation capabilities (Wang et al., 2023; Jagerman et al., 2023; Shao et al., 2023). This makes them particularly effective in situations where labeled data is scarce or unavailable. One such approach, HyDE (Hypothetical Document Embeddings)(Gao et al., 2022), employs zero-shot prompts to guide an instruction-following language model to generate hypothetical documents, effectively narrowing the semantic gap between the query and the target document. Similarly, Query2doc(Wang et al., 2023) uses few-shot prompting of LLMs to generate pseudo-documents, which are then used to expand the original query. However, applying these methods to medical information retrieval presents three critical challenges:
(1) LLMs lack the specialized medical knowledge necessary to generate highly relevant hypothetical documents. Although LLMs are trained on vast datasets drawn from a wide array of general-purpose sources, they are often insufficiently equipped with domain-specific knowledge, particularly in fields like medicine. This limitation can lead to the generation of irrelevant or misleading hypothetical documents, as LLMs may produce hallucinations—erroneous or fabricated information—due to their limited understanding of complex medical concepts.
(2) General text embedding models are inadequate for representing medical queries and documents effectively. These models are typically designed for multi-domain (e.g., medicine, law, economics) and multi-task (e.g., retrieval, question answering, classification) settings, with the aim of creating versatile retrievers capable of handling various fields, tasks, and languages. However, such generalization often fails to capture the nuanced and knowledge-intensive nature of the medical domain, where precise representation of medical terminology and context is crucial for accurate retrieval.
(3) The medical domain suffers from a scarcity of high-quality, relevance-labeled datasets. In particular, relevance-labeled medical datasets—especially in non-English languages—are extremely limited. Creating such datasets, which require the careful curation of query-document pairs, is both time-consuming and resource-intensive. The scarcity of labeled data significantly increases the cost of training and fine-tuning models to achieve high performance in medical information retrieval.
In this paper, we aim to develop an effective fully zero-shot dense retrieval system that requires no relevance-labeled data for medical information retrieval. Inspired by the concept of enhancing retrieval through hypothetical documents, we design an end-to-end framework called Self-Learning Hypothetical Document Embedding (SL-HyDE). During the inference phase, the framework first employs an LLM as a generator to produce a relevant hypothetical document in response to a medical query. A retrieval model is then employed to identify the most relevant real document based on the generated hypothetical document. In the training phase, we design a self-learning mechanism that enhances the retrieval performance of SL-HyDE without the need for labeled data. Specifically, this mechanism leverages the retrieval model’s ranking capabilities to select high-relevance hypothetical documents that align with the output of the generator, effectively injecting medical knowledge into the LLM. In turn, the generator’s ability to produce high-quality hypothetical documents provides pseudo-labeled data for the retrieval model, enabling it to efficiently encode medical texts. This interactive and complementary approach generates supervisory signals that enhance both the generation and retrieval capabilities of the system. Notably, SL-HyDE begins with unlabeled medical corpora and completes the training process through a self-learning mechanism, thereby circumventing the heavy reliance on labeled data typically required for training both large language models and retrieval systems.
To evaluate SL-HyDE’s performance in Chinese medical information retrieval, we identify a significant lack of suitable benchmarks. In the broader field of information retrieval, datasets like BEIR (Thakur et al., 2021) offer a collection of 18 datasets that assess embedding performance across a range of retrieval tasks. MTEB (Muennighoff et al., 2023) and C-MTEB (Xiao et al., 2024) provide extensive benchmarks for massive text embedding tasks (such as retrieval, ranking, and clustering) in English and Chinese, respectively. Despite these resources, Chinese medical datasets are significantly underrepresented. To bridge this gap, we develop a valuable Chinese Medical Information Retrieval Benchmark (CMIRB). CMIRB is constructed from real-world medical scenarios, including online consultations, medical examinations, and literature retrieval. It comprises five tasks and ten datasets, marking the first comprehensive and authentic evaluation benchmark for Chinese medical information retrieval. This benchmark is poised to accelerate advancements toward more robust and generalizable MIR systems in the future.
Through extensive experimentation on the CMIRB benchmark, we find that our proposed method significantly enhances retrieval performance. We validate SL-HyDE across various configurations involving three large language models as generators and three embedding models as retrievers. Notably, SL-HyDE surpasses the HyDE (Qwen2 as generator + BGE as retriever) combination by an average of 4.9% in NDCG@10 across ten datasets, and it shows a 7.2% improvement compared to using BGE alone for retrieval. These outcomes underscore the effectiveness and versatility of SL-HyDE. In summary, our contributions are as follows:
  • We propose Self-Learning Hypothetical Document Embeddings (SL-HyDE) for zero-shot medical information retrieval, eliminating the need for relevance-labeled data.
  • We develop a comprehensive Chinese Medical Information Retrieval Benchmark (CMIRB) and evaluate the performance of various retrieval models using this benchmark.
  • SL-HyDE enhances retrieval accuracy across five tasks and demonstrates generalizability and scalability with different combinations of generators and retrievers.
Figure 1. Training and inference pipeline of SL-HyDE.
2. RELATED WORK
2.1. Dense Retrieval
In recent years, with rapid advancements in deep learning and natural language processing (NLP) technologies, researchers in the field of Information Retrieval (IR) have increasingly focused on leveraging advanced text representation methods to improve retrieval system performance. Contriever (Izacard et al., 2021) leverages unsupervised contrastive learning for dense retrieval. BGE (Xiao et al., 2024) enhances Chinese general embeddings through training on large-scale text pairs. GTE (Li et al., 2023) employs multi-stage contrastive learning for multilingual applications. mE5 (Wang et al., 2024c) uses weakly supervised pretraining and fine-tuning for multilingual embeddings. BMRETRIEVER (Xu et al., 2024) specializes in English biomedical retrieval. These works demonstrate the impact of well-structured training strategies on effective retrieval across domains. Beyond embedding-based techniques, large language models (LLMs) have demonstrated exceptional performance in zero-resource retrieval scenarios, primarily due to their extensive knowledge and robust text generation capabilities. GAR (Mao et al., 2021) enriches query semantics with generated content. HyDE (Gao et al., 2022) generates hypothetical documents for the retriever, effectively narrowing the semantic gap between the query and the target document. Query2doc (Wang et al., 2023) utilizes few-shot prompts to expand queries, boosting both sparse and dense retrieval. These methods have significantly enhanced retrieval in open-domain QA.
However, retrieval through hypothetical documents generated by LLMs often yields suboptimal results when domain-specific knowledge is insufficient. To address these challenges, we propose a self-learning framework that jointly optimizes the generator and retriever without any relevance labels, thereby enhancing retrieval performance.
2.2. Information Retrieval Benchmark
To better guide the development of retrieval models, researchers have developed various datasets and benchmarks to comprehensively evaluate retrieval model performance. For instance, DuReader (He et al., 2018), a large-scale Chinese reading comprehension dataset, provides a valuable question-answer resource, significantly advancing text understanding and information retrieval research. Similarly, HotPotQA (Yang et al., 2018) introduces challenges requiring complex multi-hop reasoning, driving advancements in multi-document retrieval. BEIR(Thakur et al., 2021), a zero-shot retrieval evaluation benchmark, covers diverse retrieval tasks and offers a unified evaluation platform. MTEB (Muennighoff et al., 2023) establishes a framework for evaluating multilingual text embeddings. More recently, C-MTEB (Xiao et al., 2024) specifically addresses Chinese text embedding evaluations. Collectively, these benchmarks have propelled progress in IR technologies and provided critical resources for embedding research.
However, most benchmarks are designed for general domains, limiting their utility for specific domains such as medical retrieval. Existing medical benchmarks like CMB (Wang et al., 2024b) and CMExam (Liu et al., 2024) focus primarily on medical QA and clinical reasoning, which are not suitable for medical retrieval evaluation. To bridge this gap, we compile and adapt medical corpora specifically for IR tasks thereby improving the evaluation of medical retrievers and guiding their optimization for specialized applications.
3. METHODOLOGY
3.1. Dense Retrieval
Zero-shot document retrieval is a crucial component of the search systems. Given a user query $q$ and a document set $D=\{d_{1},...,d_{n}\}$ where n represents the number of document candidates, the goal of a retrieval model ($\mathcal{M}_{r}$) is to fetch documents that align with the user’s genuine search intent for the current query $q$. These models map an input query $q$ and a document $d$ into a pair of vectors $\textlangle v_{q},v_{d}\textrangle$, using their inner product as a similarity function $s(q,d)$:
(1) $s(q,d)=<\mathcal{M}_{r}(q),\mathcal{M}_{r}(d)>$
The retrieval models then identify the top-k documents, denoted as $\overline{D}$, which have the highest similarity scores when compared to the query $q$. Earlier studies (Karpukhin et al., 2020; Ding et al., 2020) utilized dual-tower architectures to encode queries and documents separately. However, more recent research (Xiong et al., 2020; Izacard et al., 2021; Xiao et al., 2024) indicates that using the same architecture for both query and document encoding provides greater robustness in low-resource information retrieval tasks. In this work, we utilize the $\mathcal{M}_{t}$ model to serve as both the query encoder and the document encoder.
3.2. Hypothetical Document Embedding (HyDE)
Large language models have achieved remarkable success in text generation across various natural language processing tasks, including question answering (Liu et al., 2021), text generation (Dathathri et al., 2019), and dialogue systems. Recently, there has been a growing interest in utilizing these models to generate relevant documents based on queries, thereby improving retrieval accuracy. Hypothetical Document Embeddings (HyDE) (Gao et al., 2022) decompose dense retrieval into two tasks: a generative task executed by an instruction-following language model and a document-document similarity task executed by a retrieval model. Initially, the query is fed to the generative model with the instruction to ”write a document that answers the question,” creating a hypothetical document. Specifically, a large language model such as ChatGPT or InstructGPT can serve as the generative model ($\mathcal{M}_{g}$). By deploying prompts, the user query is supplied, and the model generates the hypothetical documents:
(2) $d^{\prime}=\mathcal{M}_{g}(q,Prompt_{1})$
To better fuse the documents, they sample $N$ documents $[d^{\prime}_{1},...,d^{\prime}_{N}]$ from generated hypothetical documents. Subsequently, an unsupervised contrastive encoder (such as Contriever) is used to encode these documents into an embedding vector $v_{q}$.
(3) $v_{q}=\frac{1}{N+1}[\sum_{k=1}^{N}\mathcal{M}_{r}(d^{\prime}_{k})+\mathcal{M}_{r}(q)]$
The inner product is then computed between $\mathcal{M}_{r}(d)$ and the set of all document vectors:
(4) $s(q,d)=<v_{q},\mathcal{M}_{r}(d)>\forall d\in D$
This vector identifies a neighborhood in the corpus embedding space, from which similar real documents are retrieved based on vector similarity. This straightforward plug-and-play paradigm makes HyDE easy to deploy in real-world applications for solving retrieval tasks.
3.3. SL-HyDE
Applying HyDE to the medical domain presents two primary challenges: (1) LLMs often lack specialized medical domain knowledge, and (2) retrievers may struggle to effectively encode medical texts due to inadequate training on medical corpora. These challenges hinder the successful implementation of HyDE technology in the medical field, making it difficult to achieve significant performance improvements in retrieval tasks. A common strategy to supplement medical domain knowledge involves fine-tuning with labeled medical data (Zhang et al., 2023; Wang et al., 2024a; Tian et al., 2023; Xu et al., 2024). However, these approaches rely on high-quality, manually constructed data to adapt general models to the medical domain. Unfortunately, obtaining such high-quality labeled data in practice is particularly challenging, making the training of a medical LLM highly costly.
In this paper, we introduce a Self-Learning Hypothetical Document Embedding (SL-HyDE) mechanism designed to leverage the potential of unlabeled corpora in the medical domain. SL-HyDE consists of two main modules: a self-learning generator and a self-learning retriever. The overall framework is depicted in Figure 1.
3.4. Self-Learning Generator
A common strategy for equipping a large model with medical domain knowledge involves continued pretraining on unlabeled corpora, supervised fine-tuning on labeled data, and reinforcement learning to align the model with human preferences. However, continued pretraining typically requires substantial computational resources and extensive datasets. Consequently, supervised fine-tuning with smaller datasets has gained popularity as an efficient method for adapting large models. In this paper, we utilize the generator and retriever in SL-HyDE to create a labeled dataset based on an unsupervised corpus. The labels are entirely generated by SL-HyDE, eliminating the need for external labeled data collection.
We begin with an unlabeled corpus, specifically utilizing a medical encyclopedia dataset, Huatuo26M_encyclopedia [^1] https://huggingface.co/datasets/FreedomIntelligence/huatuo_encyclopedia_qa, as our foundational text. To construct queries, we employ a robust offline LLM, Qwen2.5222https://huggingface.co/Qwen/Qwen2.5-32B-Instruct, leveraging in-context learning (Brown, 2020; Min et al., 2022). This approach generates meaningful queries in a specific format, using prompts such as: ”Please generate a medical question based on the provided medical text. Demonstration: [DEMONSTRATION] Text: [TEXT]”.
(5) $q=LLM(d,Prompt_{2})$
After generating the query, we employ the generator to produce a hypothetical document that encapsulates the correct information based on the true target document. We intentionally avoid using the true target document as the output label because the primary role of the generator is to craft a hypothetical document that aids the retriever in locating the true target document. Expecting the generator to replicate the exact target document itself would be overly demanding and unrealistic.
(6) $\{d_{1}^{\prime},...d_{N}^{\prime}\}=\mathcal{M}_{g}(q,d,Prompt_{3})$
Considering that the hypothetical documents generated by the generator may not always be ideal for retrieving the true target document. Ideally, the optimal hypothetical document is one that enhances the ranking position of the target document during retrieval. To achieve this, we employ the retriever $\mathcal{M}_{r}$ to rank and select the optimal hypothetical document. Specifically, the generator creates $K$ hypothetical documents for a given query, and each hypothetical document is utilized by the retriever to search the corpus for the most relevant documents. We then calculate the ranking position $r_{i}$ of the true target document for $d^{\prime}_{i}$:
(7) $r_{i}=rank(d,\mathcal{M}_{r}(d^{\prime}_{i},D)),i=1,...,K$
(8) $d^{*}=d^{\prime}_{\mathrm{arg\ min^{K}_{i=1}}\ r_{i}}$
The higher the ranking, the more effective the hypothetical document. We select the generated document with the highest rank (i.e., the lowest $r_{i}$). At this stage, we have constructed a question-answer pair in the form of $(q,d^{*})$, where the query serves as the question, and the generated document acts as the answer. We then employ a supervised fine-tuning approach to train the generator using this dataset, $D_{llm}=\{(q,d^{*})|q\in Q\}$. The standard supervised fine-tuning (SFT) loss is calculated as follows:
(9) $\mathcal{L}_{\mathrm{sft}}=-\sum\nolimits_{q\in Q}\sum\nolimits_{t}\mathrm{log}\ \mathcal{M}_{g}(q^{\prime}_{t}|q^{\prime}_{\textless t},q)$
For the training data of the self-learning generator, there is no need to rely on supervision signals from labeled medical data. Instead, we utilize only unlabeled corpora, leveraging the generator and retriever within the SL-HyDE framework to create the supervision signals. More importantly, the self-learning generator trained in this manner can produce hypothetical documents optimized for retrieval ranking preferences, thereby enhancing the retriever’s performance in subsequent retrieval processes.
3.5. Self-Learning Retriever
The training of retrievers typically involves two stages: unsupervised pre-training and supervised fine-tuning. Unsupervised pre-training aims to harness large volumes of unlabeled text to learn domain knowledge, while supervised fine-tuning employs high-quality labeled datasets, often alongside hard negative mining, to further teach the model to differentiate between similar texts. For example, GTE (Li et al., 2023) utilizes $\sim$800M text pairs during the unsupervised pre-training stage and $\sim$3M pairs for fine-tuning. In this work, our goal is not to train a medical retriever with vast amounts of data but rather to adapt a general-purpose retriever to the medical domain through fine-tuning with a small amount of data.
[…]
In addition to the unlabeled medical corpus, we now have a corresponding query. This $(q,d)$ combination already forms the query-document labeled data required for retriever training. However, considering that HyDE retrieves documents by simultaneously encoding both the query and the hypothetical document to locate the target document, we propose using a $(q,d^{\prime},d)$ triplet as the labeled data for retriever training. This method narrows the gap between the training and inference stages for the retriever, ensuring the two stages operate under the same data format.
To achieve this, we utilize the fine-tuned generator $\mathcal{M}_{g}$ from the previous stage to generate hypothetical documents for all queries, constructing a labeled fine-tuning dataset $D_{emb}=\{(q,d^{\prime},d>)\}$. Following previous research (Xiao et al., 2024; Li et al., [n. d.]), we enhance the complexity of the training data through hard negative mining using the retriever, which creates challenging negative examples that closely resemble the queries. The hard negative dataset $D^{-}$ is mined from the original corpus $D$, employing the ANN-style sampling strategy in (Xiong et al., 2020):
(10) $D^{-}=ANN(\mathcal{M}_{r}(q,d^{\prime}),\mathcal{M}_{r}(d))$
In addition to the negatives mined from the corpus, we also incorporate in-batch negatives. We then apply contrastive learning loss for supervised fine-tuning of the retriever, with the objective function formulated as follows:
(11) $\mathcal{L}_{\mathrm{emb}}=\mathrm{min.}\sum_{(q,d)}-\mathrm{log}\ \frac{e^{s(q,d)/\tau}}{e^{s(q,d)/\tau}+\sum\nolimits_{B\cup D^{-}}e^{s(q,d^{-})/\tau}}$
where $\tau$ is the temperature coefficient, and $B$ represents the negative samples within the batch. The similarity score $s(q,d)$ incorporates the generated document, as described in Equation 4.
Thus far, we have developed a retriever equipped with medical domain knowledge that is coherently adapted to the characteristics of retrieval queries, incorporating hypothetical documents within the HyDE framework.
3.6. SL-HyDE vs. HyDE
Our approach, SL-HyDE, builds upon HyDE with several enhancements while retaining some similarities. First, both SL-HyDE and HyDE follow the same inference process for information retrieval tasks. Each uses a large model to generate a hypothetical document based on the query, which the retriever then employs to locate the most relevant document. Second, neither SL-HyDE nor HyDE requires labeled data, which allows for rapid deployment. HyDE is especially advantageous in real-world scenarios where efficient retrieval can be executed simply by selecting a generator and a retriever. However, for tasks needing domain-specific knowledge, such as medical information retrieval, deploying HyDE directly may not yield optimal results. One potential strategy is to fine-tune the generator and retriever separately using labeled medical data before deploying the HyDE framework. The primary challenge here in acquiring labeled data, and fine-tuning the models separately often leads to suboptimal performance.
SL-HyDE improves upon this by integrating a self-learning mechanism, transforming HyDE into a trainable end-to-end framework. This mechanism enables both the generator and the retriever to better adapt to the medical domain without relying on labeled medical data. Supervision signals for the generator’s training are derived from the retriever, and vice versa, facilitating mutual enhancement through this self-learning process. This holistic approach results in improved performance in retrieval tasks. Overall, SL-HyDE offers an efficient and convenient solution for enhancing HyDE’s performance in the medical domain, particularly when dealing with unlabeled corpora.
4. CMIRB BENCHMARK
The CMIRB benchmark is a specialized multi-task dataset designed specifically for medical information retrieval. Our collection and construction methodology is guided by the following principles: (i) Domain Specificity: The benchmark focuses exclusively on medical tasks, providing a targeted evaluation setting. (ii) Task Diversity: It includes a variety of query types and document formats to reflect real-world medical scenarios. (iii) Task Difficulty: The tasks are crafted to challenge existing models, thereby pushing the boundaries of medical retrieval research.
Figure 2. An overview of the diverse tasks and datasets in CMIRB benchmark.
4.1. Task Definition
Figure 2 provides an overview of the tasks and datasets available in CMIRB. The benchmark comprises the following five task types:
Medical Knowledge Retrieval :    Retrieve relevant medical knowledge snippets from textbooks or encyclopedias based on a given medical entity query.
Medical Consultation Retrieval :     Extract relevant doctor’s responses to online medical consultation questions posed by patients, simulating the retrieval of expert medical advice.
Medical News Retrieval :     Focus on retrieving news articles that address queries related to COVID-19.
Medical Post Retrieval :    Retrieve the content of a forum post corresponding to its title.
Medical Literature Retrieval :     Retrieve abstracts of cited references based on the title of a medical paper or find a similar paper based on the text of a given medical paper.
4.2. Data Construction
The data corpus for CMIRB is partially sourced from existing datasets and supplemented by high-quality data collected from relevant medical websites.
For the medical knowledge retrieval task, we develop three datasets: the medical exam retrieval dataset MedExam, the medical encyclopedia knowledge retrieval dataset DuBaike, and the medical disease knowledge retrieval dataset DXYDisease. For MedExam, the queries and corpus are derived from the MEDQA(MCMLE) test dataset and MedTextbook(MCMLE) (Jin et al., 2021). We concatenate the multiple-choice question and options as the query. For DuBaike, the queries are derived from the DuReader dataset (He et al., 2017), while the documents consist of medical knowledge pages from Baidu Baike [^3] https://baike.baidu.com/. These queries are generated by users while searching on Baidu Search or Baidu Zhidao. For DXYDisease, we collect question-document pairs from the online disease knowledge website DingXiangYuan [^4] https://dxy.com/. The questions consist of pre-constructed inquiries about various aspects of diseases, such as definitions and symptoms.
For the medical consultation retrieval task, we collect three online patient-doctor consultation datasets: MedicalRetrieval, CmedqaRetrieval, and DXYConsult. The MedicalRetrieval dataset is sourced from (Long et al., 2022), and the CmedqaRetrieval dataset comes from (Qiu et al., 2022). Additionally, we compile several doctor-patient dialogues from the DingXiangYuan, to create DXYConsult. Compared to the former two datasets, the query in DXYConsult includes more detailed patient information, such as symptoms, medication usage, and pending diagnosis questions.
For the medical news retrieval task, we utilize the CovidRetrieval dataset from (Qiu et al., 2022). This dataset features manually annotated high-quality questions related to COVID-19 as queries, with corresponding news articles as the documents.
In the medical post retrieval task, we curate IIYiPost by crawling posts from the IIYi forum [^5] https://bbs.iiyi.com/, a platform where medical professionals discuss various medical topics. This dataset includes both post titles and the full post content.
For the medical literature retrieval task, we develop two datasets based on the CSL data (Li et al., 2022): the citation retrieval dataset CSLCite and similar literature retrieval dataset CSLRel. To construct CSLCite, we utilize journal titles from CSL as queries and extract their cited references from WanFangMedical [^6] https://med.wanfangdata.com.cn/ as positive documents. Similarly, we select the most similar paper recommended by the website as the positive documents for the CLSRel dataset.
To ensure dataset quality, we design a filtering process, as outlined in Algorithm 1. We use ChatGPT to exclude non-medical data and remove low-quality query-document pairs. For the MedExam and DuBaike datasets, beyond the initial filtering, we implement a query-document matching process to find the most similar positive document for each query, detailed in Appendix A.1. After processing, we compile a total of 10 datasets. The statistical details of these datasets are presented in Table 1. Statistics reveal that query lengths vary significantly, ranging from a single title to a full article. Similarly, document lengths span from brief physician responses to extensive medical knowledge texts. This diversity ensures that our benchmark is both comprehensive and practically meaningful.
Table 1. Statistics of datasets in CMIRB.
#Samples Avg. Word Lengths
Task Dataset #Query #Document Query Document
Medical MedExam 697 27,871 96.9 493.7
Knowledge DuBaike 318 56,441 7.6 403.3
Retrieval DXYDisease 1,255 54,021 24.3 191.1
Medical MedicalRet. 1,000 100,999 17.9 122.0
Consultation CmedqaRet. 3,999 100,001 48.4 307.7
Retrieval DXYConsult 943 12,577 170.4 370.1
News Ret. CovidRet. 949 100,001 25.9 332.4
Post Ret. IIYiPost 789 27,570 15.9 150.1
Literature CSLCite 573 36,703 21.9 269.6
Retrieval CSLRel 439 36,758 281.8 292.2
Table 2. Performance of various Retrieval models on CMIRB benchmark. All scores denote NDCG@10. The best score among 10 base retrievers on a given dataset is marked in bold.
Task Knowledge Retrieval Consulation Retrieval News Post Literature Retrieval
Dataset MedExam DuBaike DXYDis. Medical Cmedqa DXYCon. Covid IIYiPost CSLCite CSLRel Average
Text2Vec(large) 41.39 21.13 41.52 30.93 15.53 21.92 60.48 29.47 20.21 23.01 30.56
mContriever 51.50 22.25 44.34 38.50 22.71 20.04 56.01 28.11 34.59 33.95 35.20
BM25 31.95 17.89 40.12 29.33 6.83 17.78 78.90 66.95 33.74 29.97 35.35
OpenAI-Ada-002 53.48 43.12 58.72 37.92 22.36 27.69 57.21 48.60 32.97 43.40 42.55
M3E(large) 33.29 46.48 62.57 48.66 30.73 41.05 61.33 45.03 35.79 47.54 45.25
mE5(large) 53.96 53.27 72.10 51.47 28.67 41.35 75.54 63.86 42.65 37.94 52.08
piccolo(large) 43.11 45.91 70.69 59.04 41.99 47.35 85.04 65.89 44.31 44.21 54.75
GTE(large) 41.22 42.66 70.59 62.88 43.15 46.30 88.41 63.02 46.40 49.32 55.40
BGE(large) 58.61 44.26 71.71 59.60 42.57 47.73 73.33 67.13 43.27 45.79 55.40
PEG(large) 52.78 51.68 77.38 60.96 44.42 49.30 82.56 70.38 44.74 40.38 57.46
HyDE 64.39 52.73 73.98 57.27 38.52 47.11 74.32 73.07 46.16 38.68 56.62
SL-HyDE 71.49 60.96 75.34 58.58 39.07 50.13 76.95 73.81 46.78 40.71 59.38
Improve. $\uparrow$ 11.03% $\uparrow$ 15.61% $\uparrow$ 1.84% $\uparrow$ 2.29% $\uparrow$ 1.43% $\uparrow$ 6.41% $\uparrow$ 3.54% $\uparrow$ 1.01% $\uparrow$ 1.34% $\uparrow$ 5.25% $\uparrow$ 4.87%
5. EXPERIMENTS
5.1. Experimental Setup
Implementation Details    For the training medical corpus, we sample 10,000 documents from the Huatuo26_encyclopedia dataset. In our training framework, we utilize Qwen2-7B-Instruct (Yang et al., 2024) as the generator and BGE-Large-zh-v1.5 (Xiao et al., 2024) as the retriever. Unless otherwise stated, all subsequent experiments employ the Qwen+BGE setup. Model training and evaluation are conducted on up to 5 NVIDIA A100 GPUs with 40GB of memory each. During fine-tuning of the LLM, the AdamW optimizer (Loshchilov, 2017) is employed alongside a cosine-type scheduler. Training is executed for 1 epoch with a learning rate of 1e-5 and a batch size of 2. The warmup steps for the LLM are set to 200, and the lora rank is configured at 8. For retriever training, we use the AdamW optimizer with an initial learning rate of 1e-5, applying a linear decay schedule. The batch size per GPU is set at 4, and the maximum input sequence length is limited to 512. The model is trained for 1 epoch. We set the temperature to 0.02 and mined 7 hard negatives for training.
Evaluation Settings   For simplicity, we employ the LLM to generate a single hypothetical document for each query. The retrieval model embeds all queries, hypothetical documents, and corpus documents, with similarity scores calculated using cosine similarity. Documents in the corpus are ranked for each query based on these scores, and nDCG@10 is used as the primary evaluation metric to assess retrieval effectiveness.
Baseline Models   To comprehensively evaluate CMIRB, we compare our model against several popular retrieval models. These include lexical retriever BM25 (Robertson et al., 2009); dense retrieval models such as Text2Vec-Large-Chinese [^7] https://github.com/chmlyrical/text2vec-large-chinese, PEG (Wu et al., 2023), BGE-Large-zh-v1.5 (Xiao et al., 2024), GTE-Large-zh (Li et al., [n. d.]), and Piccolo-Large-zh [^8] https://github.com/timczm/piccolo-large-zh; multilingual retrievers like mContriever (masmarco) (Izacard et al., 2021), M3E-Large [^9] https://github.com/wangyuxinwhy/uniem, mE5 (multilingual-e5-large) (Wang et al., 2024c); and OpenAI’s retriever, text-embedding-ada-002101010https://platform.openai.com/docs/guides/embeddings.
5.2. Main Results
The experimental results for various retrieval models, including SL-HyDE, on the CMIRB benchmark are presented in Table 2. We make the following observations.
(1) BM25 remains highly competitive in specific retrieval scenarios. As a lexical retriever, it ranks documents based on TF-IDF matching scores calculated between queries and documents. Despite underperforming on the overall CMIRB benchmark, it displays strong results in tasks like medical news retrieval (78.9 vs. 73.33 for BGE) and medical post retrieval (66.95 vs. 67.13 for BGE). This can be attributed to the higher keyword overlap in these datasets, which aligns with BM25’s strengths.
(2) No single retrieval model achieves optimal performance across all tasks. PEG and GTE each deliver the best performance on four datasets, while BGE and mE5 each excel in achieving the top results on one dataset. Dense models with better performance often utilize contrastive learning, pretraining on large-scale unlabeled data followed by fine-tuning on labeled data. Variations in training data distribution influence model effectiveness across different datasets, suggesting the need for specialized approaches.
(3) SL-HyDE consistently outperformed HyDE across all ten datasets. While HyDE shows slight overall improvements over BGE, it excels in medical knowledge retrieval but underperforms in medical consultation tasks. This discrepancy could be due to LLM’s stronger handling of encyclopedia-type knowledge compared to the nuanced domain of patient-doctor consultations. In contrast, SL-HyDE achieved improvements over HyDE in all tasks, owing to its self-learning mechanism, which effectively enhances medical knowledge integration within both the generator and the retriever.
5.3. Performance Analysis
The core components of SL-HyDE are the generative LLM and the retrieval model. In this section, we explore the effects of varying these components.
Effect of Different LLMs   In Table 3, we present SL-HyDE’s performance with alternative LLMs fine-tuned, such as ChatGLM3-6B (Team et al., 2024) and Llama2-7b-Chat (Touvron et al., 2023), while keeping BGE as the retriever.
Both models demonstrate performance improvements under the SL-HyDE framework compared to HyDE. For instance, we observe a 4.65% improvement with ChatGLM3 and an 8.23% improvement with the Llama2 model. However, for Llama2, HyDE shows a slight decline compared to BGE. This is likely due to the fact that the pseudo-documents generated by the English-based Llama2 contained English content, which the downstream BGE retriever struggled to encode effectively. After fine-tuning, SL-HyDE improves by approximately 8%, attributed to both the reduction of English content and the enhanced retriever’s ability to encode medical knowledge, illustrating SL-HyDE’s adaptability across different LLMs.
Table 3. Performance of different generators.
Task Know. Consu. News Post Literature Avg.(All)
ChatGLM3 as Generator + BGE as Retriever
HyDE 62.43 46.43 73.89 70.88 44.46 56.02
SL-HyDE 66.26 48.55 76.78 72.29 46.40 58.63
Improve. $\uparrow$ 6.14% $\uparrow$ 4.57% $\uparrow$ 3.91% $\uparrow$ 1.99% $\uparrow$ 4.36% $\uparrow$ 4.65%
Llama2 as Generator + BGE as Retriever
HyDE 55.74 40.62 72.90 72.22 45.30 52.48
SL-HyDE 63.66 45.44 77.17 71.99 45.75 56.80
Improve. $\uparrow$ 14.21% $\uparrow$ 11.87% $\uparrow$ 5.86% $\downarrow$ 0.32% $\uparrow$ 0.99% $\uparrow$ 8.23%
Effect with Different Retrievers   We consider fine-tuning the other two retrieval models: PEG which achieves optimal performance on CMIRB, and a multilingual retrieval mE5.
In Table 4, we observe that the standard HyDE method offers some improvement over using only the retriever, but the overall performance is significantly enhanced with the application of SL-HyDE. For example, the top-performing PEG model on the CMIRB benchmark improved from 57.46% to 60.97%, representing a substantial increase in retrieval tasks. This underscores SL-HyDE’s ability to boost retrieval performance across various retriever models.
Table 4. Performance of different retrievers.
Task Know. Consu. News Post Literature Avg.(All)
Qwen2 as Generator + mE5 as Retriever
HyDE 65.77 43.15 75.92 68.15 38.58 54.80
SL-HyDE 68.60 44.83 77.59 66.81 42.33 56.94
Improve. $\uparrow$ 4.31% $\uparrow$ 3.90% $\uparrow$ 2.20% $\downarrow$ 1.97% $\uparrow$ 9.72% $\uparrow$ 3.90%
Qwen2 as Generator + PEG as Retriever
HyDE 66.03 49.73 80.49 72.51 38.87 57.80
SL-HyDE 69.96 50.97 80.89 75.93 45.03 60.97
Improve. $\uparrow$ 5.96% $\uparrow$ 2.50% $\uparrow$ 0.50% $\uparrow$ 4.72% $\uparrow$ 15.86% $\uparrow$ 5.48%
How to Use Hypothetical Documents   In this section, we test several methods for incorporating hypothetical documents. SL-HyDE: This method encodes the original query and the hypothetical documents separately, then applies mean pooling to obtain the final query vector. SL-HyDE w/K-D: This approach generates five hypothetical documents. SL-HyDE w/D: Only the hypothetical document is used as the query for retrieval. SL-HyDE w/con: The original query and the hypothetical document are concatenated into a single string to form a new query.
Table 5 shows that the combination of the original query and hypothetical documents is optimal. Sole reliance on hypothetical documents significantly reduces performance, especially in medical consultation tasks, where original queries contain critical information. The string concatenation method introduces some performance degradation, indicating that the generated documents may contain noise at the string level, whereas average pooling effectively mitigates it. Generating multiple hypothetical documents increases coverage and improves performance across tasks. However, it often leads to a K-fold increase in inference time.
Ablation Study    To further analyze the gains brought by the internal architecture of the SL-HyDE model, we conduct two sets of ablation experiments. (1) A variant, SL-HyDE w/o BGE-FT, which uses the fine-tuned LLM as the generator and the raw BGE as the retriever; (2) A variant, SL-HyDE w/o Qwen-FT, which utilizes raw LLM as the generator and the fine-tuned BGE as the retriever.
Table 6 demonstrates that fine-tuning both components substantially enhances performance, validating the efficacy of the self-learning mechanism. Notably, fine-tuning the retriever yields greater gains, suggesting that BGE benefits significantly from domain-specific adaptation. However, our approach does not rely on labeled data and still achieves significant retrieval performance improvements.
Table 5. Performance of different fusing strategies.
Task Know. Consu. News Post Literature Avg.(All)
SL-HyDE 69.26 49.26 76.95 73.81 43.75 59.38
    w/ D. 68.00 41.86 71.94 68.02 37.36 54.43
    w/ con. 69.04 45.51 73.38 69.53 44.81 57.62
    w/ K-D. 69.30 50.17 77.38 74.55 45.42 60.12
Table 6. Performance of different variants.
Task Know. Consu. News Post Literature Avg.(All)
HyDE 63.70 47.63 74.32 73.07 42.42 56.62
SL-HyDE 69.26 49.26 76.95 73.81 43.75 59.38
     w/o BGE-FT 64.32 47.95 74.87 72.91 43.24 57.11
    w/o Qwen-FT 68.75 48.85 76.63 74.52 43.11 58.77
6. CONCLUSIONS
In this paper, we introduce an automated framework for zero-shot medical information retrieval, named SL-HyDE, which operates without the need for relevance labels. Utilizing an unlabeled medical corpus, we employ a self-learning, end-to-end training framework where the retriever guides the generator’s training, and the generator, in turn, enhances the retriever. This process integrates medical knowledge to create hypothetical documents that are more effective in retrieving target documents. Furthermore, we present a comprehensive Chinese medical information retrieval benchmark, evaluating mainstream retrieval models against this new standard. Experimental findings demonstrate that SL-HyDE consistently improves retrieval accuracy over HyDE across ten datasets. Additionally, SL-HyDE shows strong adaptability and scalability, effectively enhancing retrieval performance across various combinations of generators and retrievers.
Appendix A CMIRB DATASETS
A.1. Data Process
We curated a substantial dataset from various medical resources, as presented in Table 7, which details the source distribution and data volume. Our data preprocessing pipeline, depicted in Algorithm 1, employs prompt templates outlined in Tables 11 and 12.
Initially, we use ChatGPT [^11] https://openai.com/chatgpt to perform medical relevance detection on the texts, eliminating non-medical content (lines 3-8). Subsequently, ChatGPT assesses query-document relevance, filtering out low-relevance examples (lines 27-33). Our relevance assessment considers semantic alignment and the practical significance of data samples for their respective tasks, as highlighted in prompt 12.
For the MedExam and DuBaike datasets, the direct query-document signal isn’t initially provided. Both queries and documents in the MedExam dataset originate from Work (Jin et al., 2021), where 100 randomly selected questions have corpus documents containing evidence sufficient to answer them, verified manually by the authors. In the DuBaike dataset, queries from Baidu Search and Baidu Zhidao often match the content distribution of Baidu Baike. These factors allow us to design a query-matching algorithm to locate the valuable document.
We leverage ChatGPT’s capabilities to identify the most relevant documents. Starting with a query, we use the BM25 to retrieve the top 20 relevant documents, which GPT then ranks to identify the top 3 most relevant. Ideally, these documents should be semantically related and provide sufficient answers or evidence for the query. Therefore, ChatGPT extracts document segments as evidence details for the query.
To verify the sufficiency of this evidence, GPT generates an answer to the query based on the extracted evidence fragment. A self-verification step follows: if the GPT-generated answer aligns with the document, the document is deemed a positive match for the query. For MedExam, where queries are multiple-choice questions, we verify model answers against correct ones. For DuBaike, queries are medical knowledge questions, and answers are encyclopedic. GPT scores the generated and reference answers for consistency in expressing the same medical knowledge. This detailed process is outlined in lines 10-26.
Through this iterative loop of self-ranking, evidence searching, answering, and verification, combined with ChatGPT’s advanced knowledge capabilities, we ensure high-quality, highly relevant query-document pairs. Post-processing statistical details of these datasets are displayed in Table 1.
Algorithm 1 Data Preprocessing Pipeline
1:Input: Query set $Q$, Document set $D$, A large language model $\mathrm{LLM}$ (e.g., ChatGPT)
2:Output: High-quality, highly relevant query-document pair collection
3:// Step 1: Filter out medically irrelevant queries and documents
4:for each query $q\in Q$, $d\in D$ do
5:     $med_{score}\leftarrow\mathrm{LLM}.\text{med\_score}(q/d)$
6:     if $med_{score}<threshold$ then
7:         Remove $q/d$
8:     end if
9:end for
10:// Step 2: Matching positive query-document pair
11:if query-document matching is required then
12:     for each query $q\in Q$ do
13:         // Retrieve top-k relevant documents
14:         $d_{1},d_{2},\ldots,d_{M}\leftarrow\text{BM25}(q,D)$
15:         $d_{1},d_{2},\ldots,d_{k}\leftarrow\mathrm{LLM}.\text{reranking}(q;d_{1},d_{2},\ldots,d_{M})$
16:         // Extract evidence snippets
17:         $e_{1},e_{2},\ldots,e_{k}\leftarrow\mathrm{LLM}.\text{extract\_evidence}(q,d_{1}|d_{2}|\ldots|d_{k})$
18:         // Generate answers
19:         $a_{1},a_{2},\ldots,a_{k}\leftarrow\mathrm{LLM}.\text{answer}(q,e_{1}|e_{2}|\ldots|e_{k})$
20:         for each document $d_{i}$ do
21:              if $\mathrm{LLM}.\text{validate}(a_{i},d_{i})$ is correct then
22:                  Store $(q,d_{i})$ as a matched pair
23:              end if
24:         end for
25:     end for
26:end if
27:// Step 3: Filter out pseudo-relevant match pairs
28:for each matched pair $(q,d)$ do
29:     $rel_{score}\leftarrow\mathrm{LLM}.\text{filter\_score}(q,d)$
30:     if $rel_{score}<threshold$ then
31:         Remove $(q,d)$
32:     end if
33:end for
Table 7. Dataset Collection Sources and Quantity Statistics.
A.2. Data Example
The datasets we constructed encompass various real-world medical scenarios, with examples from 10 different datasets illustrated in Figure 3. Queries can take the form of a medical paper title, a patient’s symptom description, or an exam question. Corresponding documents include abstracts of medical papers, doctor-patient diagnostic conversations, and reference materials for exam questions.
Appendix B MODELS
B.1. Baselines
To comprehensively evaluate the performance of existing retrievers on CMIRB, we selected 10 representative models, all of which have achieved strong results on the MTEB leaderboard [^12] https://huggingface.co/spaces/mteb/leaderboard.
BM25.   BM25 (Robertson et al., 2009) is a commonly used baseline retriever which uses bag-of-words and TF-IDF to perform lexical retrieval. In this paper, BM25 is implemented with Pyserini (Lin et al., 2021) using the default hyperparameters to index snippets from all corpora.
Text2Vec [^13] https://huggingface.co/GanymedeNil/text2vec-large-chinese.   It is a cosine sentence model based on a linguistically-motivated pre-trained language model (LERT).
PEG [^14] https://huggingface.co/TownsWu/PEG.    Wu et al., (Wu et al., 2023) proposes the PEG, which is trained on more than 100 million data, encompassing a wide range of domains and covering various tasks.
BGE [^15] https://huggingface.co/BAAI/bge-large-zh-v1.5.    It takes a compound recipe to train general-purpose text embedding, including, embedding-oriented pre-training, contrastive learning with sophisticated negative sampling, and instruction-based fine-tuning.
GTE [^16] https://huggingface.co/thenlper/gte-large-zh.    It presents a multi-stage contrastive learning approach to develop text embedding model that can be applied to various tasks.
Piccolo [^17] https://huggingface.co/sensenova/piccolo-large-zh.    Piccolo is a general-purpose Chinese embedding model trained using a two-stage process with weakly supervised and manually labeled text pairs.
Contriever  [^18] https://huggingface.co/facebook/mcontriever-msmarco.    It is a multilingual dense retriever with contrastive learning, which fine-tunes the pre-trained mContriever model on MS MARCO dataset.
M3E  [^19] https://huggingface.co/moka-ai/m3e-large.    M3E (Moka Massive Mixed Embedding) is a bilingual text embedding model trained on over 22 million Chinese sentence pairs, supporting tasks like cross-lingual text similarity and retrieval.
Table 8. Evaluation Prompts for LLMs.
Q2P Prompt
Please generate a medical content paragraph to answer this question.
Question: [QUESTION]
Paragraph:
T2P Prompt
Please generate a medical content paragraph based on this title.
Title: [TITLE]
Paragraph:
P2P Prompt
Please generate a similar medical paragraph for the following text.
Text: [TEXT]
Similar Paragraph:
mE5  [^20] https://huggingface.co/intfloat/multilingual-e5-large.   Multilingual E5 text embedding models that are trained with a multi-stage pipeline, involving contrastive pre-training on 1 billion multilingual text pairs, and fine-tuning on labeled datasets.
Table 9. Performance of various Retrieval models on CMIRB benchmark. All scores denote Recall@100. The best score on a given dataset is marked in bold.
Task Knowledge Retrieval Consulation Retrieval News Post Literature Retrieval
Dataset MedExam DuBaike DXYDis. Medical Cmedqa DXYCon. Covid IIYiPost CSLCite CSLRel Average
BM25 75.61 56.92 72.91 44.20 17.26 37.33 96.47 89.98 67.19 72.66 63.05
Text2Vec(large) 89.81 79.25 78.01 52.80 42.99 64.58 88.83 74.78 61.96 70.39 70.34
mContriever 93.40 86.48 84.06 61.50 53.40 62.67 84.93 70.72 72.25 84.97 75.44
mE5(large) 93.83 98.43 96.02 70.90 57.95 80.38 97.05 91.64 77.31 91.12 85.46
M3E(large) 86.08 98.43 93.55 74.00 70.61 86.96 93.26 88.97 76.09 96.58 86.45
GTE(large) 87.52 96.54 95.86 87.00 84.95 89.50 99.47 93.41 83.25 96.58 91.41
piccolo(large) 89.67 99.06 96.81 82.80 84.81 91.09 99.47 95.69 83.07 92.25 91.47
PEG(large) 95.41 98.74 98.01 83.70 84.64 89.50 98.74 96.83 81.15 92.25 91.90
BGE(large) 97.42 98.74 96.81 81.20 82.57 91.30 98.10 95.69 80.80 96.36 91.90
Table 10. Performance of different combinations of generators and retrievers on CMIRB benchmark.
Task Knowledge Retrieval Consulation Retrieval News Post Literature Retrieval
Dataset MedExam DuBaike DXYDis. Medical Cmedqa DXYCon. Covid IIYiPost CSLCite CSLRel Average
ChatGLM3 as Generator + BGE as Retriever
HyDE 61.96 54.25 71.07 56.32 37.73 45.23 73.89 70.88 45.11 43.80 56.02
SL-HyDE 67.12 59.40 72.25 57.16 38.77 49.71 76.78 72.29 45.81 46.98 58.63
Improve. $\uparrow$ 8.33% $\uparrow$ 9.49% $\uparrow$ 1.66% $\uparrow$ 1.49% $\uparrow$ 2.76% $\uparrow$ 9.90% $\uparrow$ 3.91% $\uparrow$ 1.99% $\uparrow$ 1.55% $\uparrow$ 7.26% $\uparrow$ 4.65%
Llama2 as Generator + BGE as Retriever
HyDE 53.10 45.78 68.34 53.51 31.29 37.07 72.90 72.22 44.19 46.41 52.48
SL-HyDE 64.88 56.30 69.81 54.68 36.93 44.72 77.17 71.99 44.62 46.88 56.80
Improve. $\uparrow$ 22.18% $\uparrow$ 22.98% $\uparrow$ 2.15% $\uparrow$ 2.19% $\uparrow$ 18.02% $\uparrow$ 20.64% $\uparrow$ 5.86% $\downarrow$ 0.32% $\uparrow$ 0.97% $\uparrow$ 1.01% $\uparrow$ 8.23%
Qwen2 as Generator + mE5 as Retriever
HyDE 65.18 56.35 75.77 54.31 32.02 43.12 75.92 68.15 45.66 31.50 54.80
SL-HyDE 71.36 59.50 74.95 54.68 33.95 45.87 77.59 66.81 45.65 39.01 56.94
Improve. $\uparrow$ 9.48% $\uparrow$ 5.59% $\downarrow$ 1.08% $\uparrow$ 0.68% $\uparrow$ 6.03% $\uparrow$ 6.38% $\uparrow$ 2.20% $\downarrow$ 1.97% $\downarrow$ 0.02% $\uparrow$ 23.84% $\uparrow$ 3.90%
Qwen2 as Generator + PEG as Retriever
HyDE 64.87 55.04 78.18 58.47 41.47 49.25 80.49 72.51 43.56 34.17 57.80
SL-HyDE 72.04 60.26 77.59 59.81 40.43 52.68 80.89 75.93 47.53 42.53 60.97
Improve. $\uparrow$ 11.05% $\uparrow$ 9.48% $\downarrow$ 0.75% $\uparrow$ 2.29% $\downarrow$ 2.51% $\uparrow$ 6.96% $\uparrow$ 0.50% $\uparrow$ 4.72% $\uparrow$ 9.11% $\uparrow$ 24.47% $\uparrow$ 5.48%
OpenAI-Ada-002  [^21] https://openai.com/index/new-and-improved-embedding-model/.   It is a highly efficient text embedding model that converts natural language into dense vectors for a wide range of applications, including semantic search and similarity tasks.
For the generator, we selected three highly powerful large language models.
Qwen2222222https://huggingface.co/Qwen/Qwen2-7B-Instruct.   Qwen2 is a comprehensive suite of foundational and instruction-tuned language models, encompassing a parameter range from 0.5 to 72 billion, featuring dense models and a Mixture-of-Experts model.
ChatGLM3232323https://huggingface.co/THUDM/chatglm3-6b.   ChatGLM3-6B is a next-generation conversational pre-trained model with strong performance across tasks like semantics, reasoning, and code execution, and supports complex scenarios such as tool use and function calls.
Llama2242424https://huggingface.co/meta-llama/Llama-2-7b-chat-hf.    Llama2 is an auto-regressive language model that uses an optimized transformer architecture. The tuned versions utilize supervised fine-tuning (SFT) and reinforcement learning with human feedback (RLHF) to align with human preferences for helpfulness and safety.
B.2. Evaluation Prompt
We use the C-MTEB [^25] https://github.com/FlagOpen/FlagEmbedding/tree/master/C_MTEB framework to evaluate the performance of various retrieval models on CMIRB. To ensure stability, we set the temperature coefficient to 0 when generating hypothetical documents for LLMs, producing deterministic results. For each dataset, the prompts used to generate pseudo-documents are shown in Figure 8. The IIYIPost and CSLCite datasets utilize the T2P template to prompt LLMs to generate documents based on the given title. For the CSLRel dataset, we employ the P2P template to instruct the model to produce similar text. As for the other datasets, the Q2P template is employed by the LLM to generate answers to medical questions.
Appendix C More Experiment Results
Table 9 presents the performance of 10 retrieval models on CMIRB in terms of Recall@100. In Table 10, we present a more detailed breakdown of the performance of various LLM and retriever combinations across the 10 datasets.
Table 11. Prompts for Data Processing (1).
Medical Relevance Prompt (LLM.med_score)
You will receive a question-answer pair from Baidu Search. Your task is to evaluate whether the Q&A is related to the medical field and
output the result in JSON format.
The JSON object must include the following keys:
- ”reason”: a string explaining the reason for your judgment.
- ”label”: an int, 0/1.
Please adhere to the following steps:
- If the content mentioned in the question and answer includes medical information and is related to the medical field, the label should be 1.
- If most of the content in the question and answer is unrelated to the medical field, the label should be 0.
You need to make a judgment and provide a reason.
Please output the result as required, and do not output any other content.
Here is the text:
Question: [QUESTION]
Answer: [ANSWER]
Passage Reranking Prompt (LLM.reranking)
You will be given a medical question-answer pair (Q-A) and its Top-20 most relevant documents. Your task is to select up to 3 most relevant
and independent reference documents to verify or support the given answer accurately and efficiently.
Your output is a JSON object, which must contain the following keys:
- "$doc\_id"$: a list, the document IDs you selected, ranked by importance, such as "$id_{1}"$.
- ”explanation”: a string, the reasons for your selection of these documents.
Please adhere to the following steps:
- 1. Read the medical question and its answer.
- 2. Examine the Top-20 retrieved documents to identify which documents independently contain the essential information needed to
support or verify the provided answer.
- 3. If suitable documents are found, select up to 3 documents that can independently support the answer and explain why these documents
are relevant. If none of the documents are suitable, do not select any.
The ”explanation” should be in Chinese. and your output must always be a JSON object, do not output anything else.
Now here are the medical Q-A pair and Top-20 documents:
Q-A pair: [Q-A PAIR]
Documents: [DOCUMENTS]
Evidence Extracting Prompt (LLM.extract_evidence)
You will be given a medical question, its answer and a related document. Your task is to extract evidence spans from the document that
directly or indirectly support the answer to the medical question.
Your output is a JSON object, which must contain the following keys:
- ”evidence_spans”: a list, a list of passages.
Please adhere to the following steps:
- 1. Carefully read the medical question and its answer.
- 2. Review the content of the provided document.
or verify the provided answer.
- 3. Identify and extract the passage from the document that directly supports the correct answer to the question.
- 4. If no passage in the document can directly support the correct answer or answer the question, return an empty list.
The ”explanation” should be in Chinese. and your output must always be a JSON object, do not output anything else.
Here is the medical question, its answer, and the related document
Question: [QUESTION]
Answer: [ANSWER]
Document: [DOCUMENT]
Table 12. Prompts for Data Processing (2).
Answer by Evidence Prompt (LLM.answer)
You will be given a medical exam question and one or more evidence spans that were extracted from related documents. Your task is to
provide a detailed and comprehensive answer to the question based solely on the provided evidence spans.
Your output is a JSON object, which must contain the following keys:
- ”answer”: a string, the answer you derive from the reference documents.
- ”reason”: a detailed explanation of your reasoning process leading to the answer.
Please adhere to the following steps:
- 1. Review the exam question.
- 2. Review the provided evidence spans.
- 3. Based solely on the information contained in the evidence spans, provide a detailed and comprehensive answer to the question.
- 4. If the evidence spans do not provide sufficient information to answer the question, state ”The evidence passage can not answer the
question.” in ”answer” and explain why. If you don’t know the answer, don’t guess.
You must not use any common knowledge, personal knowledge, or external information beyond the provided evidence spans.
The ”answer” and ”reason” should be in Chinese. and your output must always be a JSON object, do not output anything else.
Now here are the exam question and reference documents.
Question: [QUESTION]
Evidence Spans: [EVIDENCE SPANS]
Validate Answer Prompt (LLM.validate)
You will be given a medical question, a reference (standard) answer, and a model-generated answer. Your task is to evaluate the content
similarity between the reference answer and the model-generated answer to determine whether they are conveying the same meaning.
Your output is a JSON object, which must contain the following keys:
- ”similarity_score”: a number between 0 and 1 indicating the content similarity between the two answers.
- ”explanation”: a detailed explanation of the similarities or differences that justify your similarity score.
Please adhere to the following steps:
- 1. Carefully read the medical question.
- 2. Review the reference answer and the model-generated answer.
- 3. Compare the two answers, focusing on content similarity—whether they convey the same meaning, and lead to the same conclusion.
- 4. Provide a similarity score between 0 and 1, where 1 indicates that the answers are identical in meaning, and 0 indicates different.
- 5. Justify your score by explaining the similarities or differences between the two answers.
The ”explanation” should be in Chinese. and your output must always be a JSON object, do not output anything else.
Now here are the question, standard answer, and generated answer.
Question: [QUESTION]
Reference Answer: [REFERENCE ANSWER]
Model-generated Answer: [MODEL-GENERATED ANSWER]
Query-Document Relevance Prompt (LLM.filter_score)
You will be given a medical search query and its associated passage. Your task is to evaluate the quality of query-passage pairs intended for
use in a medical encyclopedia knowledge retrieval evaluation dataset.
Your output is a JSON object, which must contain the following keys:
- ”quality_score”: an integer, a score from 1 to 5.
- ”explanation”: a string, providing a brief rationale for the given score.
Please adhere to the following steps:
- 1. Carefully read the query to understand the user’s information need.
- 2. Review the passage to assess its relevance and targeted content in relation to the query.
- 3. Assign a quality score from 1 to 5 and explain your reasoning.
The ”explanation” should be in Chinese. and your output must always be a JSON object, do not output anything else.
Now here are the query and passage. Query: [QUERY]
Passage: [PASSAGE]
Figure 3. An overview of the diverse datasets in CMIRB benchmark.

### v2

@@
 AutoMIR: Effective Zero-Shot Medical Information Retrieval without Relevance Labels
-Abstract.
+Abstract
 Medical information retrieval (MIR) is essential for retrieving relevant medical knowledge from diverse sources, including electronic health records, scientific literature, and medical databases.
@@
 By benchmarking ten models on CMIRB, we establish a rigorous standard for evaluating medical information retrieval systems.
-Experimental results demonstrate that SL-HyDE significantly surpasses existing methods in retrieval accuracy while showcasing strong generalization and scalability across various LLM and retriever configurations.
-CMIRB data and evaluation code are publicly available at: https://github.com/CMIRB-benchmark/CMIRB
-Medical Information Retrieval, Large Language Models, Zero-shot Retrieval
-copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Retrieval models and ranking
-1. INTRODUCTION
-Medical information retrieval (MIR) (Luo et al., 2008; Goeuriot et al., 2016) focuses on retrieving relevant medical information from extensive data sources, such as electronic health records, scientific papers, and medical knowledge databases, based on specific medical queries.
-Its applications are wide-ranging, supporting doctors in clinical decision-making (Sivarajkumar et al., 2024), assisting patients in finding health information (McGowan et al., 2009), and aiding researchers in accessing relevant studies (Zheng and Yu, 2015).
-In information retrieval (IR), there are two mainstream paradigms: lexical-based sparse retrieval methods, such as BM25 (Robertson et al., 2009), and embedding-based dense retrieval approaches (Xiong et al., 2020; Ding et al., 2020; Izacard et al., 2021).
-Dense retrievers have shown strong performance, particularly in domains where large annotated datasets are available (Karpukhin et al., 2020).
-These datasets are typically created by manually matching queries with relevant documents, a process that becomes impractical when dealing with collections containing millions or billions of items.
-As a result, many retrieval tasks face the challenge of having limited or no in-domain labeled examples.
-Given these challenges, several studies (Karpukhin et al., 2020; Xiong et al., 2020) have explored transfer learning for zero-shot retrieval, where dense retrievers are trained on large, high-resource datasets and evaluated on queries from different domains.
-MS MARCO (Bajaj et al., 2016), a dataset containing a substantial number of manually annotated query-document pairs, is commonly used for this purpose.
-An alternative to transfer learning is contrastive learning (Gao et al., 2021), which follows a two-stage pretraining and fine-tuning pipeline to develop general-purpose text embedding models.
-The pretraining stage leverages weakly supervised data gathered through large-scale web crawling, while the fine-tuning stage uses high-quality text pairs derived from data mining or manual annotation (Xiao et al., 2024; Li et al., 2023).
-However, the availability of such large-scale datasets or high-quality text pairs cannot always be assumed, particularly in non-English languages or specialized domains.
-Recently, large language models (LLMs) have demonstrated exceptional performance in zero-resource retrieval scenarios, primarily due to their extensive knowledge and robust text generation capabilities (Wang et al., 2023; Jagerman et al., 2023; Shao et al., 2023).
+Experimental results demonstrate that SL-HyDE significantly surpasses HyDE in retrieval accuracy while showcasing strong generalization and scalability across various LLM and retriever configurations.
+Our code and data are publicly available at: https://github.com/ll0ruc/AutoMIR
+AutoMIR: Effective Zero-Shot Medical Information Retrieval without Relevance Labels
+Lei Li1  and Xiangxu Zhang1 and Xiao Zhou1thanks: Xiao Zhou and Zheng Liu are corresponding authors.
+and Zheng Liu211footnotemark: 1 1Gaoling School of Artificial Intelligence, Renmin University of China 2Beijing Academy of Artificial Intelligence {leil, xansar, xiaozhou}@ruc.edu.cn, zhengliu1026@gmail.com
+1 Introduction
+Medical information retrieval (MIR) Luo et al.
+(2008); Goeuriot et al.
+(2016) focuses on retrieving relevant medical information from sources like electronic health records, scientific papers, and medical knowledge databases, based on specific medical queries.
+Its applications are wide-ranging, supporting doctors in clinical decision-making Sivarajkumar et al.
+(2024), assisting patients in finding health information McGowan et al.
+(2009), and aiding researchers in accessing relevant studies Zheng and Yu (2015).
+Dense retrievers Karpukhin et al.
+(2020); Xu et al.
+(2024) have shown strong performance with large labeled datasets in information retrieval (IR).
+Several studies Xiong et al.
+(2020); Li et al.
+(2023b); Xiao et al.
+(2024) have successfully employed contrastive learning to develop general-purpose text embedding models, achieving promising results in zero-resource retrieval scenarios.
+They leverage large-scale weakly supervised data through web crawling, or high-quality text pairs derived from data mining or manual annotation.
+However, the availability of such large-scale datasets cannot always be assumed, particularly in non-English languages or specialized domains.
+Recently, large language models (LLMs) have demonstrated exceptional performance in zero-resource retrieval scenarios Wang et al.
+(2023a); Shen et al.
+(2023); Mao et al.
+(2024), primarily due to their extensive knowledge and robust text generation capabilities.
 This makes them particularly effective in situations where labeled data is scarce or unavailable.
-One such approach, HyDE (Hypothetical Document Embeddings)(Gao et al., 2022), employs zero-shot prompts to guide an instruction-following language model to generate hypothetical documents, effectively narrowing the semantic gap between the query and the target document.
-Similarly, Query2doc(Wang et al., 2023) uses few-shot prompting of LLMs to generate pseudo-documents, which are then used to expand the original query.
-However, applying these methods to medical information retrieval presents three critical challenges:
-(1) LLMs lack the specialized medical knowledge necessary to generate highly relevant hypothetical documents.
+One such approach, HyDE Gao et al.
+(2022), employs zero-shot prompts to guide an instruction-following language model to generate hypothetical documents, effectively narrowing the semantic gap between the query and the target document.
+Similarly, Query2docWang et al.
+(2023a) uses few-shot prompting of LLMs to generate pseudo-documents, which are then used to expand the original query.
+However, applying these methods to medical information retrieval presents three critical challenges: (1) LLMs lack the specialized medical knowledge necessary to generate highly relevant hypothetical documents.
 Although LLMs are trained on vast datasets drawn from a wide array of general-purpose sources, they are often insufficiently equipped with domain-specific knowledge, particularly in fields like medicine.
-This limitation can lead to the generation of irrelevant or misleading hypothetical documents, as LLMs may produce hallucinations—erroneous or fabricated information—due to their limited understanding of complex medical concepts.
 (2) General text embedding models are inadequate for representing medical queries and documents effectively.
-These models are typically designed for multi-domain (e.g., medicine, law, economics) and multi-task (e.g., retrieval, question answering, classification) settings, with the aim of creating versatile retrievers capable of handling various fields, tasks, and languages.
-However, such generalization often fails to capture the nuanced and knowledge-intensive nature of the medical domain, where precise representation of medical terminology and context is crucial for accurate retrieval.
+These versatile retrievers are typically designed for multi-domain and multi-task settings, failing to capture the nuanced and knowledge-intensive nature of the medical domain.
 (3) The medical domain suffers from a scarcity of high-quality, relevance-labeled datasets.
-In particular, relevance-labeled medical datasets—especially in non-English languages—are extremely limited.
-Creating such datasets, which require the careful curation of query-document pairs, is both time-consuming and resource-intensive.
-The scarcity of labeled data significantly increases the cost of training and fine-tuning models to achieve high performance in medical information retrieval.
-In this paper, we aim to develop an effective fully zero-shot dense retrieval system that requires no relevance-labeled data for medical information retrieval.
-Inspired by the concept of enhancing retrieval through hypothetical documents, we design an end-to-end framework called Self-Learning Hypothetical Document Embedding (SL-HyDE).
-During the inference phase, the framework first employs an LLM as a generator to produce a relevant hypothetical document in response to a medical query.
-A retrieval model is then employed to identify the most relevant real document based on the generated hypothetical document.
+The scarcity of labeled data significantly increases the cost of training and fine-tuning these models to achieve high performance.
+To address these issues, we propose Self-Learning Hypothetical Document Embedding (SL-HyDE), an effective fully zero-shot dense retrieval system requiring no relevance-labeled data for medical information retrieval.
+During the inference phase, SL-HyDE first employs an LLM as the generator to produce a relevant hypothetical document in response to a medical query.
+A retrieval model is then employed to pinpoint the most relevant target document from the candidates based on the generated hypothetical document.
 In the training phase, we design a self-learning mechanism that enhances the retrieval performance of SL-HyDE without the need for labeled data.
-Specifically, this mechanism leverages the retrieval model’s ranking capabilities to select high-relevance hypothetical documents that align with the output of the generator, effectively injecting medical knowledge into the LLM.
-In turn, the generator’s ability to produce high-quality hypothetical documents provides pseudo-labeled data for the retrieval model, enabling it to efficiently encode medical texts.
+Specifically, this mechanism leverages the retrieval model’s ranking capabilities to select high-relevance hypothetical documents that align with the output of the generator (LLM), simultaneously injecting medical knowledge into the LLM.
+In turn, the generator’s ability to produce high-quality hypothetical documents provides pseudo-labeled data for the training of retrieval model, enabling it to efficiently encode medical texts.
 This interactive and complementary approach generates supervisory signals that enhance both the generation and retrieval capabilities of the system.
-Notably, SL-HyDE begins with unlabeled medical corpora and completes the training process through a self-learning mechanism, thereby circumventing the heavy reliance on labeled data typically required for training both large language models and retrieval systems.
-To evaluate SL-HyDE’s performance in Chinese medical information retrieval, we identify a significant lack of suitable benchmarks.
-In the broader field of information retrieval, datasets like BEIR (Thakur et al., 2021) offer a collection of 18 datasets that assess embedding performance across a range of retrieval tasks.
-MTEB (Muennighoff et al., 2023) and C-MTEB (Xiao et al., 2024) provide extensive benchmarks for massive text embedding tasks (such as retrieval, ranking, and clustering) in English and Chinese, respectively.
-Despite these resources, Chinese medical datasets are significantly underrepresented.
-To bridge this gap, we develop a valuable Chinese Medical Information Retrieval Benchmark (CMIRB).
+Notably, SL-HyDE begins with unlabeled medical corpora and completes the training process through a self-learning mechanism, thereby circumventing the heavy reliance on labeled data typically required for training both large language models and text embedding models.
+To evaluate SL-HyDE’s performance in Chinese medical information retrieval, we develop a valuable Chinese Medical Information Retrieval Benchmark (CMIRB).
 CMIRB is constructed from real-world medical scenarios, including online consultations, medical examinations, and literature retrieval.
@@
 In summary, our contributions are as follows:
-  • We propose Self-Learning Hypothetical Document Embeddings (SL-HyDE) for zero-shot medical information retrieval, eliminating the need for relevance-labeled data.
-  • We develop a comprehensive Chinese Medical Information Retrieval Benchmark (CMIRB) and evaluate the performance of various retrieval models using this benchmark.
+  • We propose Self-Learning Hypothetical Document Embeddings for zero-shot medical information retrieval, eliminating the need for relevance-labeled data.
+  • We are the first to develop a comprehensive Chinese Medical Information Retrieval Benchmark and evaluate the performance of various text embedding models on it.
   • SL-HyDE enhances retrieval accuracy across five tasks and demonstrates generalizability and scalability with different combinations of generators and retrievers.
-Figure 1. Training and inference pipeline of SL-HyDE.
-2. RELATED WORK
-2.1. Dense Retrieval
-In recent years, with rapid advancements in deep learning and natural language processing (NLP) technologies, researchers in the field of Information Retrieval (IR) have increasingly focused on leveraging advanced text representation methods to improve retrieval system performance.
-Contriever (Izacard et al., 2021) leverages unsupervised contrastive learning for dense retrieval.
-BGE (Xiao et al., 2024) enhances Chinese general embeddings through training on large-scale text pairs.
-GTE (Li et al., 2023) employs multi-stage contrastive learning for multilingual applications.
-mE5 (Wang et al., 2024c) uses weakly supervised pretraining and fine-tuning for multilingual embeddings.
-BMRETRIEVER (Xu et al., 2024) specializes in English biomedical retrieval.
-These works demonstrate the impact of well-structured training strategies on effective retrieval across domains.
-Beyond embedding-based techniques, large language models (LLMs) have demonstrated exceptional performance in zero-resource retrieval scenarios, primarily due to their extensive knowledge and robust text generation capabilities.
-GAR (Mao et al., 2021) enriches query semantics with generated content.
-HyDE (Gao et al., 2022) generates hypothetical documents for the retriever, effectively narrowing the semantic gap between the query and the target document.
-Query2doc (Wang et al., 2023) utilizes few-shot prompts to expand queries, boosting both sparse and dense retrieval.
-These methods have significantly enhanced retrieval in open-domain QA.
+2 Related Work
+2.1 Dense Retrieval
+Recent advancements in deep learning and natural language processing have driven improvements in information retrieval.
+Contriever Izacard et al.
+(2021) leverages unsupervised contrastive learning for dense retrieval.
+PEG Wu et al.
+(2023) and BGE Xiao et al.
+(2024) enhance Chinese general embeddings through training on large-scale text pairs.
+These works demonstrate the impact of well-structured training strategies on effective retrieval across multiple domains.
+Beyond embedding-based techniques, large language models have demonstrated exceptional performance in zero-resource retrieval scenarios.
+GAR Mao et al.
+(2021) enriches query semantics with generated content.
+HyDE Gao et al.
+(2022) generates hypothetical documents for the retriever, effectively narrowing the semantic gap between the query and the target document.
+Query2doc Wang et al.
+(2023a) utilizes few-shot prompts to expand queries, boosting both sparse and dense retrieval.
 However, retrieval through hypothetical documents generated by LLMs often yields suboptimal results when domain-specific knowledge is insufficient.
 To address these challenges, we propose a self-learning framework that jointly optimizes the generator and retriever without any relevance labels, thereby enhancing retrieval performance.
-2.2. Information Retrieval Benchmark
-To better guide the development of retrieval models, researchers have developed various datasets and benchmarks to comprehensively evaluate retrieval model performance.
-For instance, DuReader (He et al., 2018), a large-scale Chinese reading comprehension dataset, provides a valuable question-answer resource, significantly advancing text understanding and information retrieval research.
-Similarly, HotPotQA (Yang et al., 2018) introduces challenges requiring complex multi-hop reasoning, driving advancements in multi-document retrieval.
-BEIR(Thakur et al., 2021), a zero-shot retrieval evaluation benchmark, covers diverse retrieval tasks and offers a unified evaluation platform.
-MTEB (Muennighoff et al., 2023) establishes a framework for evaluating multilingual text embeddings.
-More recently, C-MTEB (Xiao et al., 2024) specifically addresses Chinese text embedding evaluations.
-Collectively, these benchmarks have propelled progress in IR technologies and provided critical resources for embedding research.
-However, most benchmarks are designed for general domains, limiting their utility for specific domains such as medical retrieval.
-Existing medical benchmarks like CMB (Wang et al., 2024b) and CMExam (Liu et al., 2024) focus primarily on medical QA and clinical reasoning, which are not suitable for medical retrieval evaluation.
-To bridge this gap, we compile and adapt medical corpora specifically for IR tasks thereby improving the evaluation of medical retrievers and guiding their optimization for specialized applications.
-3. METHODOLOGY
-3.1. Dense Retrieval
+2.2 Information Retrieval Benchmark
+To better guide the development of retrieval models, researchers have developed various datasets and benchmarks.
+For instance, DuReader He et al.
+(2018), a large-scale Chinese reading comprehension dataset, significantly advances text understanding and information retrieval research.
+BEIRThakur et al.
+(2021), a zero-shot retrieval evaluation benchmark, covers diverse retrieval tasks and offers a unified evaluation platform.
+MTEB Muennighoff et al.
+(2023) establishes a framework for evaluating multilingual text embeddings.
+More recently, C-MTEB Xiao et al.
+(2024) specifically addresses Chinese text embedding evaluations.
+However, these benchmarks are designed for general domains, limiting their utility for specific domains such as medical retrieval.
+Existing medical benchmarks like CMB Wang et al.
+(2024b) and CMExam Liu et al.
+(2024) focus primarily on medical QA and clinical reasoning, which are not suitable for medical retrieval evaluation.
+To bridge this gap, we develop the first comprehensive and realistic evaluation benchmark based on real-world medical scenarios for Chinese medical information retrieval tasks.
+3 Methodology
+3.1 Preliminary
 Zero-shot document retrieval is a crucial component of the search systems.
@@
 These models map an input query $q$ and a document $d$ into a pair of vectors $\textlangle v_{q},v_{d}\textrangle$, using their inner product as a similarity function $s(q,d)$:
-(1) $s(q,d)=<\mathcal{M}_{r}(q),\mathcal{M}_{r}(d)>$
-The retrieval models then identify the top-k documents, denoted as $\overline{D}$, which have the highest similarity scores when compared to the query $q$.
-Earlier studies (Karpukhin et al., 2020; Ding et al., 2020) utilized dual-tower architectures to encode queries and documents separately.
-However, more recent research (Xiong et al., 2020; Izacard et al., 2021; Xiao et al., 2024) indicates that using the same architecture for both query and document encoding provides greater robustness in low-resource information retrieval tasks.
-In this work, we utilize the $\mathcal{M}_{t}$ model to serve as both the query encoder and the document encoder.
-3.2. Hypothetical Document Embedding (HyDE)
-Large language models have achieved remarkable success in text generation across various natural language processing tasks, including question answering (Liu et al., 2021), text generation (Dathathri et al., 2019), and dialogue systems.
+$s(q,d)=<\mathcal{M}_{r}(q),\mathcal{M}_{r}(d)>.$ (1)
+The retrieval models then identify the top-k documents, denoted as $D_{topk}$, which have the highest similarity scores when compared to the query $q$.
+Large language models have achieved remarkable success in text generation across various natural language processing tasks, including question answering Liu et al.
+(2021) and text generation Dathathri et al.
+(2019).
 Recently, there has been a growing interest in utilizing these models to generate relevant documents based on queries, thereby improving retrieval accuracy.
-Hypothetical Document Embeddings (HyDE) (Gao et al., 2022) decompose dense retrieval into two tasks: a generative task executed by an instruction-following language model and a document-document similarity task executed by a retrieval model.
-Initially, the query is fed to the generative model with the instruction to ”write a document that answers the question,” creating a hypothetical document.
-Specifically, a large language model such as ChatGPT or InstructGPT can serve as the generative model ($\mathcal{M}_{g}$).
-By deploying prompts, the user query is supplied, and the model generates the hypothetical documents:
-(2) $d^{\prime}=\mathcal{M}_{g}(q,Prompt_{1})$
-To better fuse the documents, they sample $N$ documents $[d^{\prime}_{1},...,d^{\prime}_{N}]$ from generated hypothetical documents.
-Subsequently, an unsupervised contrastive encoder (such as Contriever) is used to encode these documents into an embedding vector $v_{q}$.
-(3) $v_{q}=\frac{1}{N+1}[\sum_{k=1}^{N}\mathcal{M}_{r}(d^{\prime}_{k})+\mathcal{M}_{r}(q)]$
-The inner product is then computed between $\mathcal{M}_{r}(d)$ and the set of all document vectors:
-(4) $s(q,d)=<v_{q},\mathcal{M}_{r}(d)>\forall d\in D$
-This vector identifies a neighborhood in the corpus embedding space, from which similar real documents are retrieved based on vector similarity.
-This straightforward plug-and-play paradigm makes HyDE easy to deploy in real-world applications for solving retrieval tasks.
-3.3. SL-HyDE
+Hypothetical Document Embeddings (HyDE) Gao et al.
+(2022) decompose dense retrieval into two tasks: a generative task executed by an instruction-following language model and a document-document similarity task executed by a retrieval model.
+Figure 1: Training and inference pipeline of SL-HyDE.
+3.2 Overview
 Applying HyDE to the medical domain presents two primary challenges: (1) LLMs often lack specialized medical domain knowledge, and (2) retrievers may struggle to effectively encode medical texts due to inadequate training on medical corpora.
 These challenges hinder the successful implementation of HyDE technology in the medical field, making it difficult to achieve significant performance improvements in retrieval tasks.
-A common strategy to supplement medical domain knowledge involves fine-tuning with labeled medical data (Zhang et al., 2023; Wang et al., 2024a; Tian et al., 2023; Xu et al., 2024).
+A common strategy to supplement medical domain knowledge involves fine-tuning with labeled medical data Zhang et al.
+(2023); Wang et al.
+(2024c); Xu et al.
+(2024).
 However, these approaches rely on high-quality, manually constructed data to adapt general models to the medical domain.
 Unfortunately, obtaining such high-quality labeled data in practice is particularly challenging, making the training of a medical LLM highly costly.
-In this paper, we introduce a Self-Learning Hypothetical Document Embedding (SL-HyDE) mechanism designed to leverage the potential of unlabeled corpora in the medical domain.
-SL-HyDE consists of two main modules: a self-learning generator and a self-learning retriever.
-The overall framework is depicted in Figure 1.
-3.4. Self-Learning Generator
-A common strategy for equipping a large model with medical domain knowledge involves continued pretraining on unlabeled corpora, supervised fine-tuning on labeled data, and reinforcement learning to align the model with human preferences.
-However, continued pretraining typically requires substantial computational resources and extensive datasets.
-Consequently, supervised fine-tuning with smaller datasets has gained popularity as an efficient method for adapting large models.
-In this paper, we utilize the generator and retriever in SL-HyDE to create a labeled dataset based on an unsupervised corpus.
-The labels are entirely generated by SL-HyDE, eliminating the need for external labeled data collection.
-We begin with an unlabeled corpus, specifically utilizing a medical encyclopedia dataset, Huatuo26M_encyclopedia [^1] https://huggingface.co/datasets/FreedomIntelligence/huatuo_encyclopedia_qa, as our foundational text.
-To construct queries, we employ a robust offline LLM, Qwen2.5222https://huggingface.co/Qwen/Qwen2.5-32B-Instruct, leveraging in-context learning (Brown, 2020; Min et al., 2022).
-This approach generates meaningful queries in a specific format, using prompts such as: ”Please generate a medical question based on the provided medical text.
-Demonstration: [DEMONSTRATION] Text: [TEXT]”.
-(5) $q=LLM(d,Prompt_{2})$
-After generating the query, we employ the generator to produce a hypothetical document that encapsulates the correct information based on the true target document.
-We intentionally avoid using the true target document as the output label because the primary role of the generator is to craft a hypothetical document that aids the retriever in locating the true target document.
+In this paper, we introduce a self-learning hypothetical document embedding mechanism designed to leverage the potential of unlabeled medical corpora.
+The labels are entirely generated by the generator and retriever in SL-HyDE, eliminating the need for external labeled data collection.
+Figure 1 presents the overall framework.
+3.3 SL-HyDE Training
+Self-Learning Generator.
+An unlabeled medical corpus, such as Huatuo26M Li et al.
+(2023a), serves as the foundational resource for domain-specific content.
+To construct queries, we employ a robust offline LLM, Qwen2.5-32B-Instruct Team (2024), leveraging in-context learning Brown (2020).
+With a well-designed prompt, the model effectively generates medically grounded and context-aware queries:
+$q=\mathrm{LLM}(d,\mathrm{prompt}).$ (2)
+To facilitate retrieval, the raw generator creates a hypothetical document that distills the relevant information from the true target document.
+Concretely, we provide both the query and the corresponding target document as input to the generator, along with a carefully designed prompt to guide the generation of the pseudo-document.
+$d^{\prime}=\mathcal{M}_{g}(q,d,\mathrm{prompt}).$ (3)
+Notably, we intentionally avoid using the true target document as the output label because the generator’s primary role is to craft a hypothetical document that aids the retriever in locating it.
 Expecting the generator to replicate the exact target document itself would be overly demanding and unrealistic.
-(6) $\{d_{1}^{\prime},...d_{N}^{\prime}\}=\mathcal{M}_{g}(q,d,Prompt_{3})$
-Considering that the hypothetical documents generated by the generator may not always be ideal for retrieving the true target document.
-Ideally, the optimal hypothetical document is one that enhances the ranking position of the target document during retrieval.
-To achieve this, we employ the retriever $\mathcal{M}_{r}$ to rank and select the optimal hypothetical document.
-Specifically, the generator creates $K$ hypothetical documents for a given query, and each hypothetical document is utilized by the retriever to search the corpus for the most relevant documents.
-We then calculate the ranking position $r_{i}$ of the true target document for $d^{\prime}_{i}$:
-(7) $r_{i}=rank(d,\mathcal{M}_{r}(d^{\prime}_{i},D)),i=1,...,K$
-(8) $d^{*}=d^{\prime}_{\mathrm{arg\ min^{K}_{i=1}}\ r_{i}}$
-The higher the ranking, the more effective the hypothetical document.
-We select the generated document with the highest rank (i.e., the lowest $r_{i}$).
-At this stage, we have constructed a question-answer pair in the form of $(q,d^{*})$, where the query serves as the question, and the generated document acts as the answer.
-We then employ a supervised fine-tuning approach to train the generator using this dataset, $D_{llm}=\{(q,d^{*})|q\in Q\}$.
-The standard supervised fine-tuning (SFT) loss is calculated as follows:
-(9) $\mathcal{L}_{\mathrm{sft}}=-\sum\nolimits_{q\in Q}\sum\nolimits_{t}\mathrm{log}\ \mathcal{M}_{g}(q^{\prime}_{t}|q^{\prime}_{\textless t},q)$
-For the training data of the self-learning generator, there is no need to rely on supervision signals from labeled medical data.
-Instead, we utilize only unlabeled corpora, leveraging the generator and retriever within the SL-HyDE framework to create the supervision signals.
-More importantly, the self-learning generator trained in this manner can produce hypothetical documents optimized for retrieval ranking preferences, thereby enhancing the retriever’s performance in subsequent retrieval processes.
-3.5. Self-Learning Retriever
-The training of retrievers typically involves two stages: unsupervised pre-training and supervised fine-tuning.
-Unsupervised pre-training aims to harness large volumes of unlabeled text to learn domain knowledge, while supervised fine-tuning employs high-quality labeled datasets, often alongside hard negative mining, to further teach the model to differentiate between similar texts.
-For example, GTE (Li et al., 2023) utilizes $\sim$800M text pairs during the unsupervised pre-training stage and $\sim$3M pairs for fine-tuning.
-In this work, our goal is not to train a medical retriever with vast amounts of data but rather to adapt a general-purpose retriever to the medical domain through fine-tuning with a small amount of data.
-[…]
-In addition to the unlabeled medical corpus, we now have a corresponding query.
-This $(q,d)$ combination already forms the query-document labeled data required for retriever training.
-However, considering that HyDE retrieves documents by simultaneously encoding both the query and the hypothetical document to locate the target document, we propose using a $(q,d^{\prime},d)$ triplet as the labeled data for retriever training.
-This method narrows the gap between the training and inference stages for the retriever, ensuring the two stages operate under the same data format.
-To achieve this, we utilize the fine-tuned generator $\mathcal{M}_{g}$ from the previous stage to generate hypothetical documents for all queries, constructing a labeled fine-tuning dataset $D_{emb}=\{(q,d^{\prime},d>)\}$.
-Following previous research (Xiao et al., 2024; Li et al., [n.
-d.]), we enhance the complexity of the training data through hard negative mining using the retriever, which creates challenging negative examples that closely resemble the queries.
-The hard negative dataset $D^{-}$ is mined from the original corpus $D$, employing the ANN-style sampling strategy in (Xiong et al., 2020):
-(10) $D^{-}=ANN(\mathcal{M}_{r}(q,d^{\prime}),\mathcal{M}_{r}(d))$
+Given that not all hypothetical documents generated by the generator are equally effective for retrieval, we leverage the retriever $\mathcal{M}_{r}$ to select the most optimal one.
+Specifically, the generator $\mathcal{M}_{g}$ creates $L$ hypothetical documents for a given query.
+Each hypothetical document $d^{\prime}_{i}$ is used to retrieve documents from the corpus, and we record the rank $r_{i}$ of the true target document $d$.
+The pseudo-document with the highest retrieval quality (the lowest $r_{i}$) is selected:
+$r_{i}=\mathrm{rank}(d,\mathrm{sort}(s(d^{\prime}_{i},D)),i=1,...,L,$ (4)
+$i^{*}=\mathrm{arg\ min^{L}_{i=1}}\ r_{i},\ d^{*}=d^{\prime}_{i^{*}}.$ (5)
+This process yields a collection of question-answer pairs in the form of $(q,d^{*})$, functions as the question and the generated document as the corresponding answer.
+The generator is subsequently trained via supervised fine-tuning on the resulting dataset $D_{llm}=\{(q,d^{*})|q\in Q\}$.
+The standard supervised fine-tuning (SFT) loss is computed as:
+$\mathcal{L}_{\mathrm{slg}}=-\sum\nolimits_{q\in Q}\sum\nolimits_{t}\mathrm{log}\ \mathcal{M}_{g}(d^{\prime}_{t}|d^{\prime}_{\textless t},q).$ (6)
+Interestingly, the self-learning generator is trained without relying on supervision signals from labeled medical data.
+Instead, it is based on unlabeled corpora and employs the generator’s text generation alongside the retriever’s ranking function to construct high-quality question-answer pairs tailored for hypothetical document generation.
+Self-Learning Retriever.
+Given a passage $d$ from the corpus $D$ and its corresponding query $q$, the pair $(q,d)$ naturally forms the labeled query-document data required for retriever fine-tuning.
+However, since SL-HyDE retrieves the target document by encoding both the query and a generated hypothetical document when inference, we explore a triplet $(q,d^{\prime};d)$ as the labeled data for retriever training.
+This approach effectively aligns the training data format with that of the inference stage, thereby enhancing consistency and bridging the gap between training and deployment.
+To achieve this, we utilize the fine-tuned generator $\mathcal{M}_{g}^{t}$ from the previous stage to generate hypothetical documents for all queries, constructing a labeled fine-tuning dataset $D_{emb}=\{(q,d^{\prime};d)|q\in Q\}$.
+Following previous research Li et al.
+(2023b); Xiao et al.
+(2024), we further increase the training data complexity through hard negative mining.
+Specifically, a retriever is used to identify difficult negative samples from the original corpus $D$ through an ANN-based sampling strategy Xiong et al.
+(2020), resulting in a hard negative dataset:
+$D^{-}=\mathrm{ANN}(\mathcal{M}_{r}(q,d^{\prime}),\mathcal{M}_{r}(D)).$ (7)
 In addition to the negatives mined from the corpus, we also incorporate in-batch negatives.
-We then apply contrastive learning loss for supervised fine-tuning of the retriever, with the objective function formulated as follows:
-(11) $\mathcal{L}_{\mathrm{emb}}=\mathrm{min.}\sum_{(q,d)}-\mathrm{log}\ \frac{e^{s(q,d)/\tau}}{e^{s(q,d)/\tau}+\sum\nolimits_{B\cup D^{-}}e^{s(q,d^{-})/\tau}}$
-where $\tau$ is the temperature coefficient, and $B$ represents the negative samples within the batch. The similarity score $s(q,d)$ incorporates the generated document, as described in Equation 4.
-Thus far, we have developed a retriever equipped with medical domain knowledge that is coherently adapted to the characteristics of retrieval queries, incorporating hypothetical documents within the HyDE framework.
-3.6. SL-HyDE vs. HyDE
-Our approach, SL-HyDE, builds upon HyDE with several enhancements while retaining some similarities.
-First, both SL-HyDE and HyDE follow the same inference process for information retrieval tasks.
-Each uses a large model to generate a hypothetical document based on the query, which the retriever then employs to locate the most relevant document.
-Second, neither SL-HyDE nor HyDE requires labeled data, which allows for rapid deployment.
-HyDE is especially advantageous in real-world scenarios where efficient retrieval can be executed simply by selecting a generator and a retriever.
-However, for tasks needing domain-specific knowledge, such as medical information retrieval, deploying HyDE directly may not yield optimal results.
-One potential strategy is to fine-tune the generator and retriever separately using labeled medical data before deploying the HyDE framework.
-The primary challenge here in acquiring labeled data, and fine-tuning the models separately often leads to suboptimal performance.
-SL-HyDE improves upon this by integrating a self-learning mechanism, transforming HyDE into a trainable end-to-end framework.
-This mechanism enables both the generator and the retriever to better adapt to the medical domain without relying on labeled medical data.
-Supervision signals for the generator’s training are derived from the retriever, and vice versa, facilitating mutual enhancement through this self-learning process.
-This holistic approach results in improved performance in retrieval tasks.
-Overall, SL-HyDE offers an efficient and convenient solution for enhancing HyDE’s performance in the medical domain, particularly when dealing with unlabeled corpora.
-4. CMIRB BENCHMARK
+Contrastive learning loss is then applied for the supervised fine-tuning of the retriever, with the objective function formulated as follows:
+$\mathcal{L}_{\mathrm{slr}}=\mathrm{min.}\sum_{(q,d)}-\mathrm{log}\ \frac{e^{s(q,d)/\tau}}{e^{s(q,d)/\tau}+\sum\nolimits_{B\cup D^{-}}e^{s(q,d^{-})/\tau}},$ (8)
+where $\tau$ is the temperature coefficient, and $B$ represents the negative samples in a batch. The score $s(q,d)$ incorporates the generated document, as described in Equation 1.
+At this stage, we can obtain a retriever equipped with medical domain knowledge that is coherently adapted to the characteristics of retrieval queries, incorporating hypothetical documents.
+In SL-HyDE, the generator and retriever are trained separately in a sequential manner, allowing each component to be optimized with the most appropriate supervision signal available at its respective training phase.
+3.4 SL-HyDE Inference
+As illustrated in Figure 1, the inference stage of SL-HyDE introduces a hypothesis generation step prior to conventional retrieval.
+Specifically, the input query $q$ is first rewritten by a fine-tuned generator $\mathcal{M}_{g}^{t}$ to produce a pseudo-document $d^{\prime}$, as defined by the following equation:
+$d^{\prime}=\mathcal{M}_{g}^{t}(q,\mathrm{prompt}).$ (9)
+The prompt is a manually designed instruction tailored to the requirements of each task. Detailed formulations of the prompts used in our experiments are provided in Appendix A.2.
+To better fuse the documents, we sample $N$ documents from the hypothetical documents.
+Subsequently, a tuned retriever $\mathcal{M}_{r}^{t}$ is used to encode these documents into an embedding vector $v_{q}$:
+$v_{q}=\frac{1}{N+1}[\sum_{k=1}^{N}\mathcal{M}_{r}^{t}(d^{\prime}_{k})+\mathcal{M}_{r}^{t}(q)].$ (10)
+Then, the inner product is computed between $v_{q}$ and all document vectors:
+$s(q,d)=<v_{q},\mathcal{M}_{r}^{t}(d)>,\forall d\in D.$ (11)
+This vector identifies a neighborhood in the corpus embedding space, from which similar real documents are retrieved based on vector similarity.
+4 CMIRB Benchmark
+4.1 Overview
+Figure 2: An overview of CMIRB.
 The CMIRB benchmark is a specialized multi-task dataset designed specifically for medical information retrieval.
-Our collection and construction methodology is guided by the following principles: (i) Domain Specificity: The benchmark focuses exclusively on medical tasks, providing a targeted evaluation setting.
-(ii) Task Diversity: It includes a variety of query types and document formats to reflect real-world medical scenarios.
-(iii) Task Difficulty: The tasks are crafted to challenge existing models, thereby pushing the boundaries of medical retrieval research.
-Figure 2. An overview of the diverse tasks and datasets in CMIRB benchmark.
-4.1. Task Definition
-Figure 2 provides an overview of the tasks and datasets available in CMIRB. The benchmark comprises the following five task types:
-Medical Knowledge Retrieval :    Retrieve relevant medical knowledge snippets from textbooks or encyclopedias based on a given medical entity query.
-Medical Consultation Retrieval :     Extract relevant doctor’s responses to online medical consultation questions posed by patients, simulating the retrieval of expert medical advice.
-Medical News Retrieval :     Focus on retrieving news articles that address queries related to COVID-19.
-Medical Post Retrieval :    Retrieve the content of a forum post corresponding to its title.
-Medical Literature Retrieval :     Retrieve abstracts of cited references based on the title of a medical paper or find a similar paper based on the text of a given medical paper.
-4.2. Data Construction
-The data corpus for CMIRB is partially sourced from existing datasets and supplemented by high-quality data collected from relevant medical websites.
-For the medical knowledge retrieval task, we develop three datasets: the medical exam retrieval dataset MedExam, the medical encyclopedia knowledge retrieval dataset DuBaike, and the medical disease knowledge retrieval dataset DXYDisease.
-For MedExam, the queries and corpus are derived from the MEDQA(MCMLE) test dataset and MedTextbook(MCMLE) (Jin et al., 2021).
-We concatenate the multiple-choice question and options as the query.
-For DuBaike, the queries are derived from the DuReader dataset (He et al., 2017), while the documents consist of medical knowledge pages from Baidu Baike [^3] https://baike.baidu.com/.
-These queries are generated by users while searching on Baidu Search or Baidu Zhidao.
-For DXYDisease, we collect question-document pairs from the online disease knowledge website DingXiangYuan [^4] https://dxy.com/.
-The questions consist of pre-constructed inquiries about various aspects of diseases, such as definitions and symptoms.
-For the medical consultation retrieval task, we collect three online patient-doctor consultation datasets: MedicalRetrieval, CmedqaRetrieval, and DXYConsult.
-The MedicalRetrieval dataset is sourced from (Long et al., 2022), and the CmedqaRetrieval dataset comes from (Qiu et al., 2022).
-Additionally, we compile several doctor-patient dialogues from the DingXiangYuan, to create DXYConsult.
-Compared to the former two datasets, the query in DXYConsult includes more detailed patient information, such as symptoms, medication usage, and pending diagnosis questions.
-For the medical news retrieval task, we utilize the CovidRetrieval dataset from (Qiu et al., 2022).
-This dataset features manually annotated high-quality questions related to COVID-19 as queries, with corresponding news articles as the documents.
-In the medical post retrieval task, we curate IIYiPost by crawling posts from the IIYi forum [^5] https://bbs.iiyi.com/, a platform where medical professionals discuss various medical topics.
-This dataset includes both post titles and the full post content.
-For the medical literature retrieval task, we develop two datasets based on the CSL data (Li et al., 2022): the citation retrieval dataset CSLCite and similar literature retrieval dataset CSLRel.
-To construct CSLCite, we utilize journal titles from CSL as queries and extract their cited references from WanFangMedical [^6] https://med.wanfangdata.com.cn/ as positive documents.
-Similarly, we select the most similar paper recommended by the website as the positive documents for the CLSRel dataset.
-To ensure dataset quality, we design a filtering process, as outlined in Algorithm 1.
-We use ChatGPT to exclude non-medical data and remove low-quality query-document pairs.
-For the MedExam and DuBaike datasets, beyond the initial filtering, we implement a query-document matching process to find the most similar positive document for each query, detailed in Appendix A.1.
-After processing, we compile a total of 10 datasets.
-The statistical details of these datasets are presented in Table 1.
-Statistics reveal that query lengths vary significantly, ranging from a single title to a full article.
-Similarly, document lengths span from brief physician responses to extensive medical knowledge texts.
-This diversity ensures that our benchmark is both comprehensive and practically meaningful.
-Table 1. Statistics of datasets in CMIRB.
+As shown in Figure 2, it comprises five different tasks.
+Medical knowledge retrieval task: Retrieve relevant medical knowledge snippets from textbooks or encyclopedias based on a given medical query.
+Medical consultation retrieval task: Extract relevant doctor’s responses to online medical consultation questions posed by patients.
+Medical news retrieval task: Focus on retrieving news articles that address queries related to COVID-19.
+Medical post retrieval task: Retrieve the content of a forum post corresponding to its title.
+Medical literature retrieval task: Retrieve abstracts of cited references based on a medical title or find a similar paper based on the given medical paper.
+4.2 Data Construction
+The CMIRB benchmark integrates 10 datasets, including several existing resources: MedicalRetrieval Long et al.
+(2022), CmedqaRetrieval Qiu et al.
+(2022), and CovidRetrieval Qiu et al.
+(2022), covering patient-doctor consultations and COVID-19-related news retrieval.
+In addition, we construct several datasets by combining existing query resources with curated medical corpora.
+MedExam pairs questions with textbook passages from MedQA Jin et al.
+(2021).
+DuBaike uses queries from DuReaderHe et al.
+(2017) and documents collected from Baidu Baike pages [^1] https://baike.baidu.com/.
+We also curate two datasets from the medical website DingXiangYuan [^2] https://dxy.com/.
+DXYDisease focuses on structured disease-related Q&A, while DXYConsult captures richer patient-doctor dialogues that include symptom descriptions, medication history, and diagnostic queries.
+We curate IIYiPost by crawling posts from the IIYi forum [^3] https://bbs.iiyi.com/.
+Finally, CSLCite and CSLRel are constructed based on the CSL dataset Li et al.
+(2022), targeting different literature retrieval scenarios.
+CSLCite uses journal titles as queries and their cited references from WanFangMedical [^4] https://med.wanfangdata.com.cn/ as documents, while CSLRel pairs each paper with the most relevant similar paper recommended by the platform.
+To ensure quality, we apply ChatGPT to exclude non-medical content and low-quality query-document pairs.
+Additional query-document matching is performed for MedExam and DuBaike to ensure content relevance.
+Full details are provided in the Appendix B.1.
+Table 1 summarizes dataset statistics, revealing broad variability in query and document length, ranging from short titles to long passages, ensuring the benchmark’s diversity and practical relevance.
 #Samples Avg. Word Lengths
@@
 Retrieval CSLRel 439 36,758 281.8 292.2
-Table 2. Performance of various Retrieval models on CMIRB benchmark. All scores denote NDCG@10. The best score among 10 base retrievers on a given dataset is marked in bold.
+Table 1: Statistics of datasets in CMIRB.
 Task Knowledge Retrieval Consulation Retrieval News Post Literature Retrieval
@@
 PEG(large) 52.78 51.68 77.38 60.96 44.42 49.30 82.56 70.38 44.74 40.38 57.46
+BGE(large) 58.61 44.26 71.71 59.60 42.57 47.73 73.33 67.13 43.27 45.79 55.40
 HyDE 64.39 52.73 73.98 57.27 38.52 47.11 74.32 73.07 46.16 38.68 56.62
@@
 Improve. $\uparrow$ 11.03% $\uparrow$ 15.61% $\uparrow$ 1.84% $\uparrow$ 2.29% $\uparrow$ 1.43% $\uparrow$ 6.41% $\uparrow$ 3.54% $\uparrow$ 1.01% $\uparrow$ 1.34% $\uparrow$ 5.25% $\uparrow$ 4.87%
-5. EXPERIMENTS
-5.1. Experimental Setup
-Implementation Details    For the training medical corpus, we sample 10,000 documents from the Huatuo26_encyclopedia dataset.
-In our training framework, we utilize Qwen2-7B-Instruct (Yang et al., 2024) as the generator and BGE-Large-zh-v1.5 (Xiao et al., 2024) as the retriever.
-Unless otherwise stated, all subsequent experiments employ the Qwen+BGE setup.
-Model training and evaluation are conducted on up to 5 NVIDIA A100 GPUs with 40GB of memory each.
-During fine-tuning of the LLM, the AdamW optimizer (Loshchilov, 2017) is employed alongside a cosine-type scheduler.
+Table 2: Performance of various Retrieval models on nDCG@10.
+The first part shows ten base retrieval models, and the second shows retrieval models enhanced by hypothetical documents.
+denotes the result outperforms baseline models in t-test at $p<0.05$ level.
+5 Experiments
+5.1 Experimental Setup
+Implementation Details.
+We sample 10,000 documents from the Huatuo26M_encyclopedia dataset as the unlabeled corpus.
+In our training framework, we utilize Qwen2-7B-Instruct Yang et al.
+(2024) as the generator and BGE-Large-zh-v1.5 Xiao et al.
+(2024) as the retriever.
+Unless otherwise stated, all experiments are conducted under this Qwen+BGE configuration.
+Model training and evaluation are conducted on up to 5 NVIDIA A100 GPUs, each equipped with 40GB of memory.
+For fine-tuning the LLM, we employ the AdamW optimizer Loshchilov (2017) in conjunction with a cosine learning rate scheduler.
 Training is executed for 1 epoch with a learning rate of 1e-5 and a batch size of 2.
-The warmup steps for the LLM are set to 200, and the lora rank is configured at 8.
-For retriever training, we use the AdamW optimizer with an initial learning rate of 1e-5, applying a linear decay schedule.
+We set 200 warmup steps and configure the LoRA rank to 8.
+Retriever fine-tuning also uses the AdamW optimizer, with a linear decay schedule and an initial learning rate of 1e-5.
 The batch size per GPU is set at 4, and the maximum input sequence length is limited to 512.
-The model is trained for 1 epoch.
-We set the temperature to 0.02 and mined 7 hard negatives for training.
-Evaluation Settings   For simplicity, we employ the LLM to generate a single hypothetical document for each query.
+We apply a temperature of 0.02 and mine 7 hard negatives for each query to enhance training difficulty.
+Evaluation Settings.
+For simplicity, we employ the LLM to generate a single hypothetical document for each query.
 The retrieval model embeds all queries, hypothetical documents, and corpus documents, with similarity scores calculated using cosine similarity.
 Documents in the corpus are ranked for each query based on these scores, and nDCG@10 is used as the primary evaluation metric to assess retrieval effectiveness.
-Baseline Models   To comprehensively evaluate CMIRB, we compare our model against several popular retrieval models.
-These include lexical retriever BM25 (Robertson et al., 2009); dense retrieval models such as Text2Vec-Large-Chinese [^7] https://github.com/chmlyrical/text2vec-large-chinese, PEG (Wu et al., 2023), BGE-Large-zh-v1.5 (Xiao et al., 2024), GTE-Large-zh (Li et al., [n.
-d.]), and Piccolo-Large-zh [^8] https://github.com/timczm/piccolo-large-zh; multilingual retrievers like mContriever (masmarco) (Izacard et al., 2021), M3E-Large [^9] https://github.com/wangyuxinwhy/uniem, mE5 (multilingual-e5-large) (Wang et al., 2024c); and OpenAI’s retriever, text-embedding-ada-002101010https://platform.openai.com/docs/guides/embeddings.
-5.2. Main Results
+We set the temperature of LLM to 0.7 and repeat five times with different random seeds.
+Baseline Models.
+To comprehensively evaluate CMIRB, we select several popular retrieval models.
+These include lexical retriever BM25 Robertson et al.
+(2009); dense retrieval models such as Text2Vec-Large-Chinese Xu (2023), PEG Wu et al.
+(2023), BGE-Large-zh-v1.5 Xiao et al.
+(2024), GTE-Large-zh Li et al.
+(2023b), and Piccolo-Large-zh SenseTime (2023); multilingual retrievers like mContriever (masmarco) Izacard et al.
+(2021), M3E-Large Wang et al.
+(2023b), mE5 (multilingual-e5-large) Wang et al.
+(2024a); and text-embedding-ada-002 OpenAI .
+5.2 Main Results
 The experimental results for various retrieval models, including SL-HyDE, on the CMIRB benchmark are presented in Table 2. We make the following observations.
-(1) BM25 remains highly competitive in specific retrieval scenarios.
+(1) BM25 remains highly competitive in specific medical tasks.
 As a lexical retriever, it ranks documents based on TF-IDF matching scores calculated between queries and documents.
@@
 67.13 for BGE).
-This can be attributed to the higher keyword overlap in these datasets, which aligns with BM25’s strengths.
-(2) No single retrieval model achieves optimal performance across all tasks.
+This can be attributed to the higher keyword overlap in datasets.
+(2) No single retrieval model achieves optimal performance across all ten tasks.
 PEG and GTE each deliver the best performance on four datasets, while BGE and mE5 each excel in achieving the top results on one dataset.
@@
 This discrepancy could be due to LLM’s stronger handling of encyclopedia-type knowledge compared to the nuanced domain of patient-doctor consultations.
-In contrast, SL-HyDE achieved improvements over HyDE in all tasks, owing to its self-learning mechanism, which effectively enhances medical knowledge integration within both the generator and the retriever.
-5.3. Performance Analysis
-The core components of SL-HyDE are the generative LLM and the retrieval model. In this section, we explore the effects of varying these components.
-Effect of Different LLMs   In Table 3, we present SL-HyDE’s performance with alternative LLMs fine-tuned, such as ChatGLM3-6B (Team et al., 2024) and Llama2-7b-Chat (Touvron et al., 2023), while keeping BGE as the retriever.
-Both models demonstrate performance improvements under the SL-HyDE framework compared to HyDE.
+In contrast, SL-HyDE achieved improvements over HyDE in all tasks, owing to its self-learning mechanism, which effectively enhances medical knowledge integration within both the generator and the retriever, while also aligning the outputs of the two models.
+5.3 Performance Analysis
+Effect of Different Generators.
+In Table 3, we present SL-HyDE’s performance with alternative fine-tuned LLMs as the generator, such as ChatGLM3-6B Team et al.
+(2024) and Llama2-7b-Chat Touvron et al.
+(2023).
+Both models demonstrate performance improvements under SL-HyDE compared to HyDE.
 For instance, we observe a 4.65% improvement with ChatGLM3 and an 8.23% improvement with the Llama2 model.
@@
 This is likely due to the fact that the pseudo-documents generated by the English-based Llama2 contained English content, which the downstream BGE retriever struggled to encode effectively.
-After fine-tuning, SL-HyDE improves by approximately 8%, attributed to both the reduction of English content and the enhanced retriever’s ability to encode medical knowledge, illustrating SL-HyDE’s adaptability across different LLMs.
-Table 3. Performance of different generators.
+After fine-tuning, SL-HyDE improves by approximately 8%, attributed to both the reduction of English content and the enhanced retriever’s ability to encode medical knowledge, illustrating SL-HyDE’s adaptability.
 Task Know. Consu. News Post Literature Avg.(All)
@@
 Improve. $\uparrow$ 14.21% $\uparrow$ 11.87% $\uparrow$ 5.86% $\downarrow$ 0.32% $\uparrow$ 0.99% $\uparrow$ 8.23%
-Effect with Different Retrievers   We consider fine-tuning the other two retrieval models: PEG which achieves optimal performance on CMIRB, and a multilingual retrieval mE5.
+Table 3: Performance of different generators.
+Effect of Different Retrievers. We consider fine-tuning the other two retrieval models: PEG which achieves optimal performance on CMIRB, and a multilingual retriever mE5.
 In Table 4, we observe that the standard HyDE method offers some improvement over using only the retriever, but the overall performance is significantly enhanced with the application of SL-HyDE.
@@
 This underscores SL-HyDE’s ability to boost retrieval performance across various retriever models.
-Table 4. Performance of different retrievers.
 Task Know. Consu. News Post Literature Avg.(All)
@@
 Improve. $\uparrow$ 5.96% $\uparrow$ 2.50% $\uparrow$ 0.50% $\uparrow$ 4.72% $\uparrow$ 15.86% $\uparrow$ 5.48%
-How to Use Hypothetical Documents   In this section, we test several methods for incorporating hypothetical documents.
+Table 4: Performance of different retrievers.
+Task Know. Consu. News Post Literature Avg.(All)
+SL-HyDE 69.26 49.26 76.95 73.81 43.75 59.38
+    w/ D. 68.00 41.86 71.94 68.02 37.36 54.43
+    w/ con. 69.04 45.51 73.38 69.53 44.81 57.62
+    w/ K-D. 69.30 50.17 77.38 74.55 45.42 60.12
+Table 5: Performance of different fusing strategies.
+Effect of Different Fusing Strategies.
+In this section, we test several methods for incorporating hypothetical documents.
 SL-HyDE: This method encodes the original query and the hypothetical documents separately, then applies mean pooling to obtain the final query vector.
-SL-HyDE w/K-D: This approach generates five hypothetical documents.
 SL-HyDE w/ D: Only the hypothetical document is used as the query for retrieval.
 SL-HyDE w/ con: The original query and the hypothetical document are concatenated into a single string to form a new query.
+SL-HyDE w/ K-D: This approach generates five documents.
 Table 5 shows that the combination of the original query and hypothetical documents is optimal.
@@
 However, it often leads to a K-fold increase in inference time.
-Ablation Study    To further analyze the gains brought by the internal architecture of the SL-HyDE model, we conduct two sets of ablation experiments.
-(1) A variant, SL-HyDE w/o BGE-FT, which uses the fine-tuned LLM as the generator and the raw BGE as the retriever; (2) A variant, SL-HyDE w/o Qwen-FT, which utilizes raw LLM as the generator and the fine-tuned BGE as the retriever.
+Therefore, we need to balance efficiency and accuracy to select the number of hypothetical documents.
+5.4 Ablation Study
+To further analyze the gains brought by the internal architecture of SL-HyDE, we conduct two sets of ablation experiments: (1) SL-HyDE w/o BGE-FT, which uses the fine-tuned LLM as the generator and the raw BGE as the retriever; (2) SL-HyDE w/o Qwen-FT, which utilizes raw LLM as the generator and the fine-tuned BGE as the retriever.
 Table 6 demonstrates that fine-tuning both components substantially enhances performance, validating the efficacy of the self-learning mechanism.
 Notably, fine-tuning the retriever yields greater gains, suggesting that BGE benefits significantly from domain-specific adaptation.
-However, our approach does not rely on labeled data and still achieves significant retrieval performance improvements.
-Table 5. Performance of different fusing strategies.
-Task Know. Consu. News Post Literature Avg.(All)
-SL-HyDE 69.26 49.26 76.95 73.81 43.75 59.38
-    w/ D. 68.00 41.86 71.94 68.02 37.36 54.43
-    w/ con. 69.04 45.51 73.38 69.53 44.81 57.62
-    w/ K-D. 69.30 50.17 77.38 74.55 45.42 60.12
-Table 6. Performance of different variants.
+However, our approach fine-tunes both the retriever and the generator, boosting their performance between the two to enhance retrieval tasks.
 Task Know. Consu. News Post Literature Avg.(All)
@@
     w/o Qwen-FT 68.75 48.85 76.63 74.52 43.11 58.77
-6. CONCLUSIONS
+Table 6: Performance of different variants.
+5.5 Case Study
+To intuitively show how the SL-HyDE makes a difference in the hypothetical documents and retrieval performance, we present examples in Table 7 to compare the hypothetical document generated by HyDE and SL-HyDE.
+The query is How to treat a hernia?.
+While HyDE generates a general document discussing conservative and surgical treatments, it lacks specificity for different patient groups.
+In contrast, SL-HyDE produces a document mentioning hernia belts for infants and elderly patients, closely matching the target document’s details.
+This improved relevance led to a higher retrieval ranking ($2nd$ vs.
+$10th$), demonstrating how more precise hypothetical documents enhance retrieval performance.
+Query: How to treat a hernia?
+Target Doc: Inguinal Hernia Treatment Plan. For conventional treatment,
+a 1-year-old infant can use a hernia belt for compression. As the muscles
+gradually strengthen, there may be a possibility of spontaneous recovery.
+For elderly and frail individuals a hernia belt can be worn, but for
+other patients, surgery is generally recommended…
+HyDE: Hernia is a common disease caused by a weak area in the
+abdominal wall, Treatment usually includes conservative and surgical
+methods. For most patients, especially young and healthy individuals,
+surgery is the preferred option… (Rank: 10)
+SL-HyDE: Hernia is a common condition that typically occurs… For
+infants,… the use of a hernia belt to apply localized pressure can help
+alleviate symptoms and promote the development of the abdominal
+muscles,… For elderly or frail patients, or those with severe underlying
+conditions,… wearing a hernia belt can help manage symptoms and
+reduce the risk of the hernia progressing further… (Rank: 2)
+Table 7: The case study comparing with baseline.
+6 Conclusions
 In this paper, we introduce an automated framework for zero-shot medical information retrieval, named SL-HyDE, which operates without the need for relevance labels.
@@
 Additionally, SL-HyDE shows strong adaptability and scalability, effectively enhancing retrieval performance across various combinations of generators and retrievers.
-Appendix A CMIRB DATASETS
-A.1. Data Process
-We curated a substantial dataset from various medical resources, as presented in Table 7, which details the source distribution and data volume.
-Our data preprocessing pipeline, depicted in Algorithm 1, employs prompt templates outlined in Tables 11 and 12.
-Initially, we use ChatGPT [^11] https://openai.com/chatgpt to perform medical relevance detection on the texts, eliminating non-medical content (lines 3-8).
-Subsequently, ChatGPT assesses query-document relevance, filtering out low-relevance examples (lines 27-33).
-Our relevance assessment considers semantic alignment and the practical significance of data samples for their respective tasks, as highlighted in prompt 12.
-For the MedExam and DuBaike datasets, the direct query-document signal isn’t initially provided.
-Both queries and documents in the MedExam dataset originate from Work (Jin et al., 2021), where 100 randomly selected questions have corpus documents containing evidence sufficient to answer them, verified manually by the authors.
-In the DuBaike dataset, queries from Baidu Search and Baidu Zhidao often match the content distribution of Baidu Baike.
-These factors allow us to design a query-matching algorithm to locate the valuable document.
-We leverage ChatGPT’s capabilities to identify the most relevant documents.
-Starting with a query, we use the BM25 to retrieve the top 20 relevant documents, which GPT then ranks to identify the top 3 most relevant.
-Ideally, these documents should be semantically related and provide sufficient answers or evidence for the query.
-Therefore, ChatGPT extracts document segments as evidence details for the query.
-To verify the sufficiency of this evidence, GPT generates an answer to the query based on the extracted evidence fragment.
-A self-verification step follows: if the GPT-generated answer aligns with the document, the document is deemed a positive match for the query.
-For MedExam, where queries are multiple-choice questions, we verify model answers against correct ones.
-For DuBaike, queries are medical knowledge questions, and answers are encyclopedic.
-GPT scores the generated and reference answers for consistency in expressing the same medical knowledge.
-This detailed process is outlined in lines 10-26.
-Through this iterative loop of self-ranking, evidence searching, answering, and verification, combined with ChatGPT’s advanced knowledge capabilities, we ensure high-quality, highly relevant query-document pairs.
-Post-processing statistical details of these datasets are displayed in Table 1.
-Algorithm 1 Data Preprocessing Pipeline
-1:Input: Query set $Q$, Document set $D$, A large language model $\mathrm{LLM}$ (e.g., ChatGPT)
-2:Output: High-quality, highly relevant query-document pair collection
-3:// Step 1: Filter out medically irrelevant queries and documents
-4:for each query $q\in Q$, $d\in D$ do
-5:     $med_{score}\leftarrow\mathrm{LLM}.\text{med\_score}(q/d)$
-6:     if $med_{score}<threshold$ then
-7:         Remove $q/d$
-8:     end if
-9:end for
-10:// Step 2: Matching positive query-document pair
-11:if query-document matching is required then
-12:     for each query $q\in Q$ do
-13:         // Retrieve top-k relevant documents
-14:         $d_{1},d_{2},\ldots,d_{M}\leftarrow\text{BM25}(q,D)$
-15:         $d_{1},d_{2},\ldots,d_{k}\leftarrow\mathrm{LLM}.\text{reranking}(q;d_{1},d_{2},\ldots,d_{M})$
-16:         // Extract evidence snippets
-17:         $e_{1},e_{2},\ldots,e_{k}\leftarrow\mathrm{LLM}.\text{extract\_evidence}(q,d_{1}|d_{2}|\ldots|d_{k})$
-18:         // Generate answers
-19:         $a_{1},a_{2},\ldots,a_{k}\leftarrow\mathrm{LLM}.\text{answer}(q,e_{1}|e_{2}|\ldots|e_{k})$
-20:         for each document $d_{i}$ do
-21:              if $\mathrm{LLM}.\text{validate}(a_{i},d_{i})$ is correct then
-22:                  Store $(q,d_{i})$ as a matched pair
-23:              end if
-24:         end for
-25:     end for
-26:end if
-27:// Step 3: Filter out pseudo-relevant match pairs
-28:for each matched pair $(q,d)$ do
-29:     $rel_{score}\leftarrow\mathrm{LLM}.\text{filter\_score}(q,d)$
-30:     if $rel_{score}<threshold$ then
-31:         Remove $(q,d)$
-32:     end if
-33:end for
-Table 7. Dataset Collection Sources and Quantity Statistics.
-A.2. Data Example
-The datasets we constructed encompass various real-world medical scenarios, with examples from 10 different datasets illustrated in Figure 3.
-Queries can take the form of a medical paper title, a patient’s symptom description, or an exam question.
-Corresponding documents include abstracts of medical papers, doctor-patient diagnostic conversations, and reference materials for exam questions.
-Appendix B MODELS
-B.1. Baselines
+In future work, we will extend SL-HyDE to other data-scarce domains to further evaluate its generalizability across different settings.
+In addition, we will explore reinforcement learning to train more capable retrievers and enhance reasoning in complex medical retrieval tasks.
+7 Limitations
+While our work effectively addresses the adaptation challenges of HyDE in low-resource scenarios, several limitations remain.
+First, our study primarily focuses on the medical domain and provides a preliminary exploration in the legal domain (see Appendix A.4), but we have not extended our investigation to other vertical domains such as economics or education.
+Second, although we experiment with three open-source LLMs, Qwen2, LlaMA2, and ChatGLM3, as generators, we do not include more recent or diverse model families such as Qwen3 or Gemini, which may exhibit different generation behaviors.
+Third, our data construction pipeline relies on LLMs for query-document matching and pseudo-relevant pair filtering.
+The effectiveness of these components depends on the model’s instruction-following ability and its sensitivity to domain-specific nuances, which may introduce hallucinations or spurious correlations.
+Appendix A Models
+A.1 Baselines
 To comprehensively evaluate the performance of existing retrievers on CMIRB, we selected 10 representative models, all of which have achieved strong results on the MTEB leaderboard [^5] https://huggingface.co/spaces/mteb/leaderboard.
-BM25.
-BM25 (Robertson et al., 2009) is a commonly used baseline retriever which uses bag-of-words and TF-IDF to perform lexical retrieval.
-In this paper, BM25 is implemented with Pyserini (Lin et al., 2021) using the default hyperparameters to index snippets from all corpora.
-Text2Vec [^13] https://huggingface.co/GanymedeNil/text2vec-large-chinese.   It is a cosine sentence model based on a linguistically-motivated pre-trained language model (LERT).
-PEG [^14] https://huggingface.co/TownsWu/PEG.
-Wu et al., (Wu et al., 2023) proposes the PEG, which is trained on more than 100 million data, encompassing a wide range of domains and covering various tasks.
-BGE [^15] https://huggingface.co/BAAI/bge-large-zh-v1.5.
+For details regarding the retrievers and large reasoning models evaluated throughout the paper, please refer to Table 8.
+BM25 Robertson et al.
+(2009).
+BM25 is a commonly used baseline retriever which uses bag-of-words and TF-IDF to perform lexical retrieval.
+In this paper, BM25 is implemented with Pyserini Lin et al.
+(2021) using the default hyperparameters to index snippets from all corpora.
+Text2Vec Xu (2023).   It is a cosine sentence model based on a linguistically-motivated pre-trained language model (LERT).
+PEG Wu et al. (2023).    Wu et al., Wu et al. (2023) proposes the PEG, which is trained on more than 100 million data, encompassing a wide range of domains and covering various tasks.
+BGE Xiao et al.
+(2024).
 It takes a compound recipe to train general-purpose text embedding, including, embedding-oriented pre-training, contrastive learning with sophisticated negative sampling, and instruction-based fine-tuning.
-GTE [^16] https://huggingface.co/thenlper/gte-large-zh.    It presents a multi-stage contrastive learning approach to develop text embedding model that can be applied to various tasks.
-Piccolo [^17] https://huggingface.co/sensenova/piccolo-large-zh.
-Piccolo is a general-purpose Chinese embedding model trained using a two-stage process with weakly supervised and manually labeled text pairs.
-Contriever  [^18] https://huggingface.co/facebook/mcontriever-msmarco.
-It is a multilingual dense retriever with contrastive learning, which fine-tunes the pre-trained mContriever model on MS MARCO dataset.
-M3E  [^19] https://huggingface.co/moka-ai/m3e-large.
+GTE Li et al. (2023b).    It presents a multi-stage contrastive learning approach to develop text embedding model that can be applied to various tasks.
+Piccolo SenseTime (2023).    Piccolo is a general-purpose Chinese embedding model trained using a two-stage process with weakly supervised and manually labeled text pairs.
+Contriever  Izacard et al. (2021).    It is a multilingual dense retriever with contrastive learning, which fine-tunes the pre-trained mContriever model on MS MARCO dataset.
+M3E Wang et al.
+(2023b).
 M3E (Moka Massive Mixed Embedding) is a bilingual text embedding model trained on over 22 million Chinese sentence pairs, supporting tasks like cross-lingual text similarity and retrieval.
-Table 8. Evaluation Prompts for LLMs.
 Q2P Prompt
-Please generate a medical content paragraph to answer this question.
-Question: [QUESTION]
+Please generate a medical content paragraph to answer
+this question.
+Question: QUESTION
 Paragraph:
 T2P Prompt
-Please generate a medical content paragraph based on this title.
-Title: [TITLE]
+Please generate a medical content paragraph based on
+this title.
+Title: TITLE
 Paragraph:
 P2P Prompt
-Please generate a similar medical paragraph for the following text.
-Text: [TEXT]
+Please generate a similar medical paragraph for the
+following text.
+Text: TEXT
 Similar Paragraph:
-mE5  [^20] https://huggingface.co/intfloat/multilingual-e5-large.
+Table 9: Evaluation prompts for generators.
+mE5 Wang et al.
+(2024a).
 Multilingual E5 text embedding models that are trained with a multi-stage pipeline, involving contrastive pre-training on 1 billion multilingual text pairs, and fine-tuning on labeled datasets.
-Table 9. Performance of various Retrieval models on CMIRB benchmark. All scores denote Recall@100. The best score on a given dataset is marked in bold.
+OpenAI-Ada-002 OpenAI .
+It is a highly efficient text embedding model that converts natural language into dense vectors for a wide range of applications, including semantic search and similarity tasks.
+For the generator, we selected three highly powerful large language models.
+Qwen2 Yang et al.
+(2024).
+Qwen2 is a comprehensive suite of foundational and instruction-tuned language models, encompassing a parameter range from 0.5 to 72 billion, featuring dense models and a Mixture-of-Experts model.
+ChatGLM3 Team et al.
+(2024).
+ChatGLM3-6B is a next-generation conversational pre-trained model with strong performance across tasks like semantics, reasoning, and code execution, and supports complex scenarios such as tool use and function calls.
+Llama2 Touvron et al.
+(2023).
+Llama2 is an auto-regressive language model that uses an optimized transformer architecture.
+The tuned versions utilize supervised fine-tuning (SFT) and reinforcement learning with human feedback (RLHF) to align with human preferences for helpfulness and safety.
+A.2 Evaluation Settings
+We use the C-MTEB [^6] C-MTEB framework to evaluate the performance of various retrieval models on CMIRB.
+To ensure stability, we set the temperature of LLM to 0.7 and repeat five times with different random seeds.
+For each dataset, the prompts used to generate pseudo-documents are shown in Figure 9.
+The IIYIPost and CSLCite datasets utilize the T2P template to prompt LLMs to generate documents based on the given title.
+For the CSLRel dataset, we employ the P2P template to instruct the model to produce similar text.
+As for the other datasets, the Q2P template is employed by the LLM to generate answers to medical questions.
+A.3 SL-HyDE vs. HyDE
+Our approach, SL-HyDE, builds upon HyDE Gao et al.
+(2022) with several enhancements while retaining some similarities.
+First, both SL-HyDE and HyDE follow the same inference process.
+Each uses a large model to generate a hypothetical document based on the query, which the retriever then employs to locate the most relevant document.
+Second, neither SL-HyDE nor HyDE requires labeled data, which allows for rapid deployment.
+HyDE is especially advantageous in real-world scenarios where efficient retrieval can be executed simply by selecting a generator and a retriever.
+However, for tasks needing domain-specific knowledge, such as medical information retrieval, deploying HyDE directly may not yield optimal results.
+One potential strategy is to fine-tune the generator and retriever separately using labeled medical data before deploying the HyDE framework.
+The primary challenge here in acquiring labeled data, and fine-tuning the models separately often leads to suboptimal performance.
+SL-HyDE improves upon this by integrating a self-learning mechanism, transforming HyDE into a trainable end-to-end framework.
+This mechanism enables both the generator and the retriever to better adapt to the medical domain.
+Supervision signals for the generator’s training are derived from the retriever, and vice versa, facilitating mutual enhancement through this self-learning process.
+This holistic approach results in improved performance in retrieval tasks.
+Overall, SL-HyDE offers an efficient and convenient solution for enhancing HyDE’s performance in the medical domain, particularly when dealing with unlabeled corpora.
+A.4 More Experiment Results
+Table 10 presents the performance of 10 retrieval models on CMIRB in terms of Recall@100.
+In Table 11, we present a more detailed breakdown of the performance of various LLM and retriever combinations across the 10 datasets.
+SL-HyDE can be easily applied to other domains that lack labeled data.
+By fine-tuning both the generator and retriever using only a small amount of unstructured domain text, it builds an effective retrieval system.
+Specifically, we apply SL-HyDE to the English legal domain.
+We sample 10k law texts from pile-of-law  [^7] https://huggingface.co/datasets/pile-of-law/pile-of-law and use Llama-2-7b-chat-hf as the generator and BGE-Large-en-V1.5 as the retriever.
+We evaluate three information retrieval datasets in the law domain from MTEB.
+The results in Table 12 shows that SL-HyDE (77.25%) significantly outperforms HyDE (75.52%) in the legal domain.
 Task Knowledge Retrieval Consulation Retrieval News Post Literature Retrieval
@@
 BGE(large) 97.42 98.74 96.81 81.20 82.57 91.30 98.10 95.69 80.80 96.36 91.90
-Table 10. Performance of different combinations of generators and retrievers on CMIRB benchmark.
+Table 10: Performance of various Retrieval models on CMIRB benchmark. All scores denote Recall@100. The best score on a given dataset is marked in bold.
 Task Knowledge Retrieval Consulation Retrieval News Post Literature Retrieval
@@
 $\uparrow$ 11.05% $\uparrow$ 9.48% $\downarrow$ 0.75% $\uparrow$ 2.29% $\downarrow$ 2.51% $\uparrow$ 6.96% $\uparrow$ 0.50% $\uparrow$ 4.72% $\uparrow$ 9.11% $\uparrow$ 24.47% $\uparrow$ 5.48%
-OpenAI-Ada-002  [^21] https://openai.com/index/new-and-improved-embedding-model/.
-It is a highly efficient text embedding model that converts natural language into dense vectors for a wide range of applications, including semantic search and similarity tasks.
-For the generator, we selected three highly powerful large language models.
-Qwen2222222https://huggingface.co/Qwen/Qwen2-7B-Instruct.
-Qwen2 is a comprehensive suite of foundational and instruction-tuned language models, encompassing a parameter range from 0.5 to 72 billion, featuring dense models and a Mixture-of-Experts model.
-ChatGLM3232323https://huggingface.co/THUDM/chatglm3-6b.
-ChatGLM3-6B is a next-generation conversational pre-trained model with strong performance across tasks like semantics, reasoning, and code execution, and supports complex scenarios such as tool use and function calls.
-Llama2242424https://huggingface.co/meta-llama/Llama-2-7b-chat-hf.
-Llama2 is an auto-regressive language model that uses an optimized transformer architecture.
-The tuned versions utilize supervised fine-tuning (SFT) and reinforcement learning with human feedback (RLHF) to align with human preferences for helpfulness and safety.
-B.2. Evaluation Prompt
-We use the C-MTEB [^25] https://github.com/FlagOpen/FlagEmbedding/tree/master/C_MTEB framework to evaluate the performance of various retrieval models on CMIRB.
-To ensure stability, we set the temperature coefficient to 0 when generating hypothetical documents for LLMs, producing deterministic results.
-For each dataset, the prompts used to generate pseudo-documents are shown in Figure 8.
-The IIYIPost and CSLCite datasets utilize the T2P template to prompt LLMs to generate documents based on the given title.
-For the CSLRel dataset, we employ the P2P template to instruct the model to produce similar text.
-As for the other datasets, the Q2P template is employed by the LLM to generate answers to medical questions.
-Appendix C More Experiment Results
-Table 9 presents the performance of 10 retrieval models on CMIRB in terms of Recall@100.
-In Table 10, we present a more detailed breakdown of the performance of various LLM and retriever combinations across the 10 datasets.
-Table 11. Prompts for Data Processing (1).
-Medical Relevance Prompt (LLM.med_score)
-You will receive a question-answer pair from Baidu Search. Your task is to evaluate whether the Q&A is related to the medical field and
-output the result in JSON format.
-The JSON object must include the following keys:
-- ”reason”: a string explaining the reason for your judgment.
-- ”label”: an int, 0/1.
-Please adhere to the following steps:
-- If the content mentioned in the question and answer includes medical information and is related to the medical field, the label should be 1.
+Table 11: Performance of different combinations of generators and retrievers on CMIRB benchmark.
+Dataset legal_ legalbench_ legalbench_ Average
+summar. contracts_qa lobbying
+BGE 59.99 73.52 91.51 75.01
+HyDE 58.95 74.82 92.78 75.52
+SL-HyDE 63.50 75.10 93.15 77.25
+Table 12: Performance of SL-HyDE in legal domain.
+Appendix B CMIRB Datasets
+Table 13: Dataset collection sources and quantity statistics.
+B.1 Data Process
+We curated a substantial dataset from various medical resources, as presented in Table 13, which details the source distribution and data volume.
+Our data preprocessing pipeline, depicted in Figuer 3 and Algorithm 1, employs prompt templates outlined in Figure 4 and Figure 5.
+Figure 3: CMIRB benchmark construction pipeline.
+Initially, we use ChatGPT [^8] https://openai.com/chatgpt to perform medical relevance detection on the texts, eliminating non-medical content (lines 3-8).
+Subsequently, ChatGPT assesses query-document relevance, filtering out low-relevance examples (lines 27-33).
+Our relevance assessment considers semantic alignment and the practical significance of data samples for their respective tasks, as highlighted in prompt LABEL:tab:data-prompts2.
+For the MedExam and DuBaike datasets, the direct query-document signal isn’t initially provided.
+Both queries and documents in the MedExam dataset originate from Work Jin et al.
+(2021), where 100 randomly selected questions have corpus documents containing evidence sufficient to
+Algorithm 1 Data Preprocessing Pipeline
+1:Input: Query set $Q$, Document set $D$, A large language model $\mathrm{LLM}$ (e.g., ChatGPT)
+2:Output: High-quality, highly relevant query-document pair collection
+3:// Step 1: Filter out medically irrelevant
+4:for each query $q\in Q$, $d\in D$ do
+5:     $med_{score}\leftarrow\mathrm{LLM}.\text{med\_score}(q/d)$
+6:     if $med_{score}<threshold$ then
+7:         Remove $q/d$
+8:     end if
+9:end for
+10:// Step 2: Matching positive pairs
+11:if query-document matching then
+12:     for each query $q\in Q$ do
+13:         // Retrieve top-k documents
+14:         $D_{k}\leftarrow\text{BM25}(q,D)$
+15:         $D_{k}\leftarrow\mathrm{LLM}.\text{reranking}(q;D_{k})$
+16:         // Extract evidence snippets
+17:         $E_{k}\leftarrow\mathrm{LLM}.\text{extract\_evidence}(q,D_{k})$
+18:         // Generate answers
+19:         $A_{k}\leftarrow\mathrm{LLM}.\text{answer}(q,E_{k})$
+20:         for each document $d_{i}$ do
+21:              if $\mathrm{LLM}.\text{validate}(a_{i},d_{i})$ then
+22:                  Store $(q,d_{i})$
+23:              end if
+24:         end for
+25:     end for
+26:end if
+27:// Step 3: Filter out pseudo-relevant pairs
+28:for each matched pair $(q,d)$ do
+29:     $rel_{score}\leftarrow\mathrm{LLM}.\text{filter\_score}(q,d)$
+30:     if $rel_{score}<threshold$ then
+31:         Remove $(q,d)$
+32:     end if
+33:end for
+answer them, verified manually by the authors.
+In the DuBaike dataset, queries from Baidu Search and Baidu Zhidao often match the content distribution of Baidu Baike.
+These factors allow us to design a query-matching algorithm to locate the valuable document.
+We leverage ChatGPT’s capabilities to identify the most relevant documents.
+Starting with a query, we use the BM25 to retrieve the top 20 relevant documents, which GPT then ranks to identify the top 3 most relevant.
+Ideally, these documents should be semantically related and provide sufficient answers or evidence for the query.
+Therefore, ChatGPT extracts document segments as evidence details for the query.
+To verify the sufficiency of this evidence, GPT generates an answer to the query based on the extracted evidence fragment.
+A self-verification step follows: if the GPT-generated answer aligns with the document, the document is deemed a positive match for the query.
+For MedExam, where queries are multiple-choice questions, we verify model answers against correct ones.
+For DuBaike, queries are medical knowledge questions, and answers are encyclopedic.
+GPT scores the generated and reference answers for consistency in expressing the same medical knowledge.
+This detailed process is outlined in lines 10-26.
+Through this iterative loop of self-ranking, evidence searching, answering, and verification, combined with ChatGPT’s advanced knowledge capabilities, we ensure high-quality, highly relevant query-document pairs.
+B.2 Data Example
+The datasets we constructed encompass various real-world medical scenarios, with examples from 10 different datasets illustrated in Table 14 and Table 15.
+Queries can take the form of a medical paper title, a patient’s symptom description, or an exam question.
+Corresponding documents include abstracts of medical papers, doctor-patient diagnostic conversations, and reference materials for exam questions.
+Medical Relevance Prompt You will receive a question-answer pair from Baidu Search.
+Your task is to evaluate whether the Q&A is related to the medical field and output the result in JSON format.
+The JSON object must include the following keys: - "reason": a string explaining the reason for your judgment.
+- "label": an int, 0/1.
+Please adhere to the following steps: - If the content mentioned in the question and answer includes medical information and is related to the medical field, the label should be 1.
 - If most of the content in the question and answer is unrelated to the medical field, the label should be 0.
@@
 Please output the result as required, and do not output any other content.
-Here is the text:
-Question: [QUESTION]
-Answer: [ANSWER]
-Passage Reranking Prompt (LLM.reranking)
-You will be given a medical question-answer pair (Q-A) and its Top-20 most relevant documents. Your task is to select up to 3 most relevant
-and independent reference documents to verify or support the given answer accurately and efficiently.
-Your output is a JSON object, which must contain the following keys:
-- "$doc\_id"$: a list, the document IDs you selected, ranked by importance, such as "$id_{1}"$.
-- ”explanation”: a string, the reasons for your selection of these documents.
-Please adhere to the following steps:
-- 1. Read the medical question and its answer.
-- 2. Examine the Top-20 retrieved documents to identify which documents independently contain the essential information needed to
-support or verify the provided answer.
-- 3. If suitable documents are found, select up to 3 documents that can independently support the answer and explain why these documents
-are relevant. If none of the documents are suitable, do not select any.
-The ”explanation” should be in Chinese. and your output must always be a JSON object, do not output anything else.
-Now here are the medical Q-A pair and Top-20 documents:
-Q-A pair: [Q-A PAIR]
-Documents: [DOCUMENTS]
-Evidence Extracting Prompt (LLM.extract_evidence)
-You will be given a medical question, its answer and a related document. Your task is to extract evidence spans from the document that
-directly or indirectly support the answer to the medical question.
-Your output is a JSON object, which must contain the following keys:
-- ”evidence_spans”: a list, a list of passages.
-Please adhere to the following steps:
-- 1. Carefully read the medical question and its answer.
-- 2. Review the content of the provided document.
-or verify the provided answer.
-- 3. Identify and extract the passage from the document that directly supports the correct answer to the question.
-- 4. If no passage in the document can directly support the correct answer or answer the question, return an empty list.
-The ”explanation” should be in Chinese. and your output must always be a JSON object, do not output anything else.
-Here is the medical question, its answer, and the related document
-Question: [QUESTION]
-Answer: [ANSWER]
-Document: [DOCUMENT]
-Table 12. Prompts for Data Processing (2).
-Answer by Evidence Prompt (LLM.answer)
-You will be given a medical exam question and one or more evidence spans that were extracted from related documents. Your task is to
-provide a detailed and comprehensive answer to the question based solely on the provided evidence spans.
-Your output is a JSON object, which must contain the following keys:
-- ”answer”: a string, the answer you derive from the reference documents.
-- ”reason”: a detailed explanation of your reasoning process leading to the answer.
-Please adhere to the following steps:
-- 1. Review the exam question.
-- 2. Review the provided evidence spans.
-- 3. Based solely on the information contained in the evidence spans, provide a detailed and comprehensive answer to the question.
-- 4. If the evidence spans do not provide sufficient information to answer the question, state ”The evidence passage can not answer the
-question.” in ”answer” and explain why. If you don’t know the answer, don’t guess.
+Here is the text: Question: [QUESTION] Answer: [ANSWER]
+Passage Reranking Prompt You will be given a medical question, a reference (standard) answer, and a model-generated answer.
+Your task is to evaluate the content similarity between the reference answer and the model-generated answer to determine whether they are conveying the same meaning.
+Your output is a JSON object, which must contain the following keys: - "similarity_score": a number between 0 and 1 indicating the content similarity between the two answers.
+- "explanation": a detailed explanation of the similarities or differences that justify your similarity score.
+Please adhere to the following steps: - 1.
+Carefully read the medical question.
+- 2.
+Review the reference answer and the model-generated answer.
+- 3.
+Compare the two answers, focusing on content similarity—whether they convey the same meaning, and lead to the same conclusion.
+- 4.
+Provide a similarity score between 0 and 1, where 1 indicates that the answers are identical in meaning, and 0 indicates different.
+- 5.
+Justify your score by explaining the similarities or differences between the two answers.
+The "explanation" should be in Chinese.
+and your output must always be a JSON object, do not output anything else.
+Now here are the question, standard answer, and generated answer.
+Question: [QUESTION] Reference Answer: [REFERENCE ANSWER] Model-generated Answer: [MODEL-GENERATED ANSWER]
+Evidence Extracting Prompt You will be given a medical question, its answer and a related document.
+Your task is to extract evidence spans from the document that directly or indirectly support the answer to the medical question.
+Your output is a JSON object, which must contain the following keys: - "evidence_spans": a list, a list of passages.
+Please adhere to the following steps: - 1.
+Carefully read the medical question and its answer.
+- 2.
+Review the content of the provided document.
+- 3.
+Identify and extract the passage from the document that directly supports the correct answer to the question.
+- 4.
+If no passage in the document can directly support the correct answer or answer the question, return an empty list.
+The "explanation" should be in Chinese.
+and your output must always be a JSON object, do not output anything else.
+Here is the medical question, its answer, and the related document Question: [QUESTION] Answer: [ANSWER] Document: [DOCUMENT]
+Figure 4: Prompt for data processing (I).
+Answer by Evidence Prompt You will be given a medical exam question and one or more evidence spans that were extracted from related documents.
+Your task is to provide a detailed and comprehensive answer to the question based solely on the provided evidence spans.
+Your output is a JSON object, which must contain the following keys: - "answer": a string, the answer you derive from the reference documents.
+- "reason": a detailed explanation of your reasoning process leading to the answer.
+Please adhere to the following steps: - 1.
+Review the exam question.
+- 2.
+Review the provided evidence spans.
+- 3.
+Based solely on the information contained in the evidence spans, provide a detailed and comprehensive answer to the question.
+- 4.
+If the evidence spans do not provide sufficient information to answer the question, state "The evidence passage can not answer the question." in "answer" and explain why.
+If you don’t know the answer, don’t guess.
 You must not use any common knowledge, personal knowledge, or external information beyond the provided evidence spans.
-The ”answer” and ”reason” should be in Chinese. and your output must always be a JSON object, do not output anything else.
+The "answer" and "reason" should be in Chinese.
+and your output must always be a JSON object, do not output anything else.
 Now here are the exam question and reference documents.
-Question: [QUESTION]
-Evidence Spans: [EVIDENCE SPANS]
-Validate Answer Prompt (LLM.validate)
-You will be given a medical question, a reference (standard) answer, and a model-generated answer. Your task is to evaluate the content
-similarity between the reference answer and the model-generated answer to determine whether they are conveying the same meaning.
-Your output is a JSON object, which must contain the following keys:
-- ”similarity_score”: a number between 0 and 1 indicating the content similarity between the two answers.
-- ”explanation”: a detailed explanation of the similarities or differences that justify your similarity score.
-Please adhere to the following steps:
-- 1. Carefully read the medical question.
-- 2. Review the reference answer and the model-generated answer.
-- 3. Compare the two answers, focusing on content similarity—whether they convey the same meaning, and lead to the same conclusion.
-- 4. Provide a similarity score between 0 and 1, where 1 indicates that the answers are identical in meaning, and 0 indicates different.
-- 5. Justify your score by explaining the similarities or differences between the two answers.
-The ”explanation” should be in Chinese. and your output must always be a JSON object, do not output anything else.
+Question: [QUESTION] Evidence Spans: [EVIDENCE SPANS]
+Validate Answer Prompt You will be given a medical question, a reference (standard) answer, and a model-generated answer.
+Your task is to evaluate the content similarity between the reference answer and the model-generated answer to determine whether they are conveying the same meaning.
+Your output is a JSON object, which must contain the following keys: - "similarity_score": a number between 0 and 1 indicating the content similarity between the two answers.
+- "explanation": a detailed explanation of the similarities or differences that justify your similarity score.
+Please adhere to the following steps: - 1.
+Carefully read the medical question.
+- 2.
+Review the reference answer and the model-generated answer.
+- 3.
+Compare the two answers, focusing on content similarity—whether they convey the same meaning, and lead to the same conclusion.
+- 4.
+Provide a similarity score between 0 and 1, where 1 indicates that the answers are identical in meaning, and 0 indicates different.
+- 5.
+Justify your score by explaining the similarities or differences between the two answers.
+The "explanation" should be in Chinese.
+and your output must always be a JSON object, do not output anything else.
 Now here are the question, standard answer, and generated answer.
-Question: [QUESTION]
-Reference Answer: [REFERENCE ANSWER]
-Model-generated Answer: [MODEL-GENERATED ANSWER]
-Query-Document Relevance Prompt (LLM.filter_score)
-You will be given a medical search query and its associated passage. Your task is to evaluate the quality of query-passage pairs intended for
-use in a medical encyclopedia knowledge retrieval evaluation dataset.
-Your output is a JSON object, which must contain the following keys:
-- ”quality_score”: an integer, a score from 1 to 5.
-- ”explanation”: a string, providing a brief rationale for the given score.
-Please adhere to the following steps:
-- 1. Carefully read the query to understand the user’s information need.
-- 2. Review the passage to assess its relevance and targeted content in relation to the query.
-- 3. Assign a quality score from 1 to 5 and explain your reasoning.
-The ”explanation” should be in Chinese. and your output must always be a JSON object, do not output anything else.
-Now here are the query and passage. Query: [QUERY]
-Passage: [PASSAGE]
-Figure 3. An overview of the diverse datasets in CMIRB benchmark.
+Question: [QUESTION] Reference Answer: [REFERENCE ANSWER] Model-generated Answer: [MODEL-GENERATED ANSWER]
+Query-Document Relevance Prompt You will be given a medical search query and its associated passage.
+Your task is to evaluate the quality of query-passage pairs intended for use in a medical encyclopedia knowledge retrieval evaluation dataset.
+Your output is a JSON object, which must contain the following keys: - "quality_score": an integer, a score from 1 to 5.
+- "explanation": a string, providing a brief rationale for the given score.
+Please adhere to the following steps: - 1.
+Carefully read the query to understand the user’s information need.
+- 2.
+Review the passage to assess its relevance and targeted content in relation to the query.
+- 3.
+Assign a quality score from 1 to 5 and explain your reasoning.
+The "explanation" should be in Chinese.
+and your output must always be a JSON object, do not output anything else.
+Now here are the query and passage.
+Query: [QUERY] Passage: [PASSAGE]
+Figure 5: Prompt for data processing (II).
+MedExam
+{CJK}UTF8gbsn Query: 问题: 胃癌最常发生的转移途径是()。选项: A:直接蔓延, B:血性转移, C:种植转移, D:淋巴转移, E:沿肠管转移.
+(EN) Question: The most common metastasis route for gastric cancer is ().
+Options: A: Direct spread, B: Hematogenous metastasis, C: Seeding metastasis, D: Lymphatic metastasis, E: Along the intestinal tract.
+{CJK}UTF8gbsn Document: 外科学 3.胃癌的扩散与转移 (2)淋巴转移:是胃癌的主要转移途径,进展期胃癌的淋巴转移率高达70%左右,侵及黏膜下层的早期胃癌淋巴转移率近20%。通常将引流胃的淋巴结分为16组,有的组还可以进一步分为若千亚组…
+(EN) Surgery 3.
+Gastric cancer dissemination and metastasis (2) Lymphatic metastasis: It is the primary route of metastasis for gastric cancer, with a lymphatic metastasis rate of about 70% in advanced gastric cancer and approximately 20% in early gastric cancer invading the submucosa.
+Lymph nodes draining the stomach are usually classified into 16 groups, with some groups further divided into several subgroups…
+DuBaike
+{CJK}UTF8gbsn Query: 强迫症的表现是什么?
+(EN) What are the manifestations of obsessive-compulsive disorder (OCD)?
+{CJK}UTF8gbsn Document: 强迫症 临床表现 多发人群焦虑症与遗传因素、个性特点、不良事件、应激因素等均有关系,尤其与患者的个性特点紧密相关,比如:过分追求完美、犹豫不决、谨小慎微、固执等,具备这些不良个性特征容易患强迫症…
+(EN) Obsessive-Compulsive Disorder Clinical Manifestations Prevalent Population Anxiety disorders are related to genetic factors, personality traits, adverse events, and stress factors, particularly closely linked to the patient’s personality traits.
+For instance, excessive perfectionism, indecisiveness, meticulousness, and stubbornness are traits that increase the risk of developing OCD…
+DXYDisease
+{CJK}UTF8gbsn Query: 维生素 A 缺乏症者需要做哪些检查来诊断?
+(EN) What tests are needed to diagnose vitamin A deficiency?
+{CJK}UTF8gbsn Document: 最准确的就是血液学检查。抽血检查血清维生素 A 的水平,对于成人来说,如果在 1.05~3.15 μmol/L,那么就表明不存在维生素 A 缺乏。 如果低于参考范围下限,那就是维生素 A 缺乏了。 …
+(EN) The most accurate test is a hematological examination.
+A blood test to check the serum vitamin A levels is conducted.
+For adults, if the levels are between 1.05 and 3.15 umol/L, it indicates that there is no vitamin A deficiency.
+If the levels are below the lower limit of the reference range, it indicates vitamin A deficiency….
+MedicalRetrieval
+{CJK}UTF8gbsn Query: 一般宝宝的肚脐眼要多久愈合?
+(EN) How long does it take for a baby’s belly button to heal?
+{CJK}UTF8gbsn Document: 你好,宝宝的肚脐一般是1-2周左右会好的,时间长的也有一个月的,不过这个时候可能会有脐茸了。
+(EN) Hello, a baby’s belly button generally heals in about 1 to 2 weeks, although it may take up to a month in some cases. During this time, there might also be umbilical granuloma.
+CmedqaRetrieval
+{CJK}UTF8gbsn Query: 甲状腺手术后多久可以干活?
+(EN) How long after thyroid surgery can one return to work?
+{CJK}UTF8gbsn Document: 皮肤的修复一般由两周左右就会不影响你的颈部活动了,至于皮下软组织以及肌肉组织的修复可能时间长一下,一般一个月后就不会有明显影响了,你就可以工作了。工作中注意不要劳累,调整好自己的情绪。
+(EN) The skin usually heals in about two weeks, and you should no longer have restrictions on neck movement.
+However, the repair of subcutaneous soft tissue and muscle tissue may take longer.
+Generally, after about a month, there should be no significant impact, and you can return to work.
+During work, be sure to avoid overexertion and manage your emotions well.
+Table 14: Data example in CMIRB (I).
+DXYConsult
+{CJK}UTF8gbsn Query: 症状及患病时长:感冒,鼻炎,失去嗅觉一周。就医及用药情况:未就医,自行服用泰诺。需要解答的问题:鼻炎,失去嗅觉怎么办
+(EN) Symptoms and Duration of Illness: Cold, rhinitis, loss of smell for one week.
+Medical Consultation and Medication: No medical consultation, self-medicated with Tylenol.
+Questions Needing Answers: What to do about rhinitis and loss of smell?
+{CJK}UTF8gbsn Document: 你好,如果近期有这种感冒的病史的话,就会导致出现嗅觉功能下降,建议在口服感冒药的技术上的话,用海盐水冲洗鼻腔,一天两次,鼻喷辅舒良或者内舒拿看看效果,如果分泌过多的话,可以口服桉柠蒎胶囊,每天三次每次一粒。
+(EN) Hello, if there has been a recent history of cold symptoms, this can lead to decreased olfactory function.
+It is recommended to use saline nasal irrigation twice a day while taking cold medicine.
+You may also try nasal sprays like Budesonide or Fluticasone to see if they help.
+If there is excessive secretion, you can take Eucalyptus and Menthol capsules, three times a day, one capsule each time.
+CovidRetrieval
+{CJK}UTF8gbsn Query: 如何对待因履行工作职责感染新冠肺炎的医务人员?
+(EN) How should healthcare workers who contract COVID-19 while fulfilling their duties be treated?
+{CJK}UTF8gbsn Document: …为进一步加强疫情防控期间医务人员防护工作,切实保障医务人员身心健康,现将有关要求通知如下:一、高度重视医务人员防护工作做好医务人员防护工作,是预防和减少医务人员感染的关键举措,…
+(EN) …To further enhance the protection of healthcare workers during the pandemic and ensure their physical and mental well-being, the following requirements are hereby notified: Pay great attention to the protection of healthcare workers Ensuring proper protection for healthcare workers is a key measure in preventing and reducing infections among them, …
+IIYiPost
+{CJK}UTF8gbsn Query: 病例讨论:静脉输入阿昔洛韦2天,出现腰痛、尿少
+(EN) Case Discussion: Two days of intravenous acyclovir, followed by lower back pain and reduced urine output
+{CJK}UTF8gbsn Document: 1.病例资料,患者,男,31岁。因静脉输入阿昔洛韦2天,出现腰痛、尿少伴恶心、呕吐6天入院。患者8天前因受凉感冒,出现咳嗽、发热(最高体温38.6℃),无明显咳痰,院外静脉给予NS500ml+青霉素钠盐800万U,vd,1次/日,…
+(EN) Case Data, Patient: Male, 31 years old.
+The patient was admitted after experiencing lower back pain and reduced urine output, accompanied by nausea and vomiting for six days following two days of intravenous acyclovir administration.
+Eight days prior, the patient had caught a cold due to exposure, presenting with a cough and fever (highest temperature of 38.6°C), without significant sputum production.
+He received intravenous administration of …
+CSLCite
+{CJK}UTF8gbsn Query: 微球在组织工程中的应用
+(EN) Application of Microspheres in Tissue Engineering
+{CJK}UTF8gbsn Document: 背景:骨组织工程骨构建中如何使生长因子持续高效发挥作用是影响成骨速度和质量的关键,现多以各种材料的微球或支架作为缓释载体,但缓释作用有待提高.目的:实验拟制备壳聚糖微球,然后复合到纳米羟基磷灰石/聚乳酸羟基乙酸支架上…
+(EN) Background: In bone tissue engineering, maintaining the sustained and efficient activity of growth factors is key to influencing the speed and quality of bone formation.
+Currently, microspheres or scaffolds made from various materials are commonly used as sustained-release carriers, but the release efficiency needs improvement.
+Objective: This experiment aims to prepare chitosan microspheres and incorporate them into a nano-hydroxyapatite/polylactic-co-glycolic acid (nHA/PLGA) scaffold, …
+CSLRel
+{CJK}UTF8gbsn Query: 高血压病的辨治及预防 高血压病可归属中医学"眩晕"、"头痛"等范畴,其起病隐匿,不易引起患者的充分重视,中后期可致心脑血管疾病、肾损害…
+(EN) Differentiation and Treatment of Hypertension and Its Prevention Hypertension can be categorized under the terms "dizziness" and "headache" in traditional Chinese medicine (TCM).
+Its onset is insidious, often not receiving enough attention from patients, …
+{CJK}UTF8gbsn Document: 辨证施治高血压 高血压病是现代医学病名,在中医归属眩晕病范畴,中医认为高血压与风、火、痰、虚有关.高血压的界定根据世界卫生组织(WHO)的标准,成人在休息状态下,收缩压持续高于140毫米汞柱…
+(EN) TCM Syndrome Differentiation and Treatment of Hypertension Hypertension is a modern medical term, categorized under dizziness in TCM.
+TCM holds that hypertension is related to wind, fire, phlegm, and deficiency …
+Table 15: Data example in CMIRB (II).